### v2

# Revisiting Non-Rotating Star Models: Classical Existence and Uniqueness Theory and Scaling Relations

Abstract
This paper presents a systematic study of the properties of non-rotating stellar models governed by the Euler-Poisson system under general equations of state, including the case of polytropic gaseous stars. We revisit and extend existence results by Auchmuty and Beals [4], adapt the uniqueness results from the quantum mechanical framework of Lieb and Yau [42] to the classical Newtonian mechanical setting. The results are also synthesized in McCann [46] but without proof. The second work we do is applying a scaling method to establish relations between solutions with different total masses. As the mass tends to zero, we analyze convergence properties of the density functions and identify precise rates for the contraction or extension of their supports.
Key words: Gaseous stars, Euler-Poisson equations, Uniqueness, Scaling methods, Vanishing mass limit, Asymptotic behavior
1 Introduction
The motion of stars and planets has long been a classical subject in celestial mechanics, astrophysics, and mathematics. Historical models such as heliocentrism, proposed by Copernicus, and later studied by Newton and others on stellar rotation, form the foundation of rotating star or planet models. Early developments by Maclaurin, Jacobi, Poincaré, and Liapounov focused on incompressible, homogeneous or nearly homogeneous bodies like the Earth. For gaseous bodies such as Jupiter, stars, or systems like galaxies, density inhomogeneity and compressibility need to be accounted for. In these contexts, such bodies can be modeled as self-gravitating fluids, governed by the Euler–Poisson equations which have been extensively analyzed under various assumptions and classes of equations of state. We consider a fluid with density $\rho(x,t)\geq 0$, velocity $v(x,t)\in\mathbb{R}^{3}$, and gravitational potential $V(x,t)\in\mathbb{R}$, where $x\in\mathbb{R}^{3}$ is position and $t\geq 0$ is time. Assuming the pressure $P(\rho)$ depends on the density only (with appropriate assumptions to be described later), the Euler–Poisson system is given by
$\begin{split}\partial_{t}\rho+\nabla\cdot\left({\rho v}\right)&=0,\\ \rho\partial_{t}v+\rho\left({v\cdot\nabla}\right)v+\nabla P(\rho)&=\rho\nabla V,\\ \Delta V&=-4\pi\rho.\end{split}$ (EP)
The first attempts to construct axisymmetric rotating solutions to the compressible Euler–Poisson equations were made by astrophysicists Milne [47] and Zeipel [59] in the early 1920s, followed by Chandrasekhar [11] and Lichtenstein [41] in 1933. Further historical accounts can be found in [13, 35].
Significant progress has been made for the single rotating star problem, where the star rotates about its axis. A special case of the rotating system is when the velocity $v$ is zero. In this case, the system is actually non-rotating, and the density $\rho$ is time-independent (stationary). Under these conditions, the Euler-Poisson equations reduce to the following equation:
$\nabla P\left(\tilde{\rho}\left(x\right)\right)-\tilde{\rho}\left(x\right)\left(\nabla V_{\tilde{\rho}}\left(x\right)\right)=0$ (EP’)
where $\tilde{\rho}$ is a density function (in general with compact support), and $V_{\tilde{\rho}}(x)={\int_{\mathbb{R}^{3}}{\frac{\tilde{\rho}(y)}{\left|{y-x}\right|}\,dy}}$.
For the non-rotating case, when the pressure satisfies a polytropic law, the solutions are known as Lane–Emden solutions. However, to obtain solutions for more general pressure laws and more general angular momentum (thus including the rotating case), two approaches have been employed. The first is variational method, as used by Auchmuty and Beals [3, 4] to establish the existence of axisymmetric equilibria for compressible fluids; see also [2, 4, 9, 14, 40]. The second is a perturbative method based on the implicit function theorem, applied near non-rotating stars [31, 32, 30, 55, 56]. Nonlinear dynamical stability of rotating stars was shown in [44, 45] via variational techniques, while stability theory for non-rotating stars can be found in [20, 34, 50].
Each method has its advantages and limitations. For example, the variational approach typically requires $\gamma>\frac{4}{3}$ when $P(\rho)=K\rho^{\gamma}$ to prevent gravitational collapse (see Remark 2.13 and Proposition 2.16 of this paper, [46, Section 2], or [4, Section 1, 6 and 8]). In contrast, perturbative methods can handle the range $\frac{6}{5}<\gamma<2$ [32], but they rely on prior knowledge of solutions in special non-rotating cases.
These methods are quite general and apply not only to single-body problems but also to multi-body systems. Variational methods were used by McCann [46] for binary stars and star-planet systems in the authors’ accompanying work [16]. Perturbative approaches appear in the work of Alonso-Orán, Kepka, and Velázquez [1] on incompressible Euler–Poisson systems with an external particle, as well as in Lichtenstein’s construction of rotating binary stars [41].
Researchers have also considered systems with more than two stellar objects, such as galaxies or general $N$-star systems. For rotating galaxies, single rotating galaxies were constructed in [53, 55] in the spirit of Lichtenstein and Heilig [30, 41]. There is extensive literature on non-rotating galaxies: see [6, 7, 26, 51] and references therein. The orbital stability of stationary solutions has seen considerable progress over the past two decades [28, 23, 24, 25, 22, 27, 36, 37, 38, 39, 51, 58], with recent work on linearly oscillating galaxies in [29]. In [10], Campos, Del Pino, and Dolbeault constructed $N$-body rotating galaxies by connecting them to relative equilibria (point particles) in $N$-body dynamics with small uniform angular velocity. They perturbed radial equilibria, and solved for the gravitational potential rather than the density, using a finite-dimensional reduction. Although natural, Jang and Seok remark in [33] that the framework in [10] is not ideal for stability analysis. They adapt Rein’s reduction method [50, 51] to extend the discussion of binary star solutions to establish existence and stability for rotating binary galaxies modeled by the Vlasov–Poisson system [33]. Furthermore, they note that for $N\geq 3$, uniformly rotating $N$-body stellar configurations lack a variational structure analogous to the binary case and are not generally expected to be stable [33].
While recent research has extended to rotating multi-body systems, a thorough understanding of the simpler, non-rotating single-body problem remains very important. In addition to the existence result established via variational methods by Auchmuty and Beals [4], Lieb and Yau [42] contributed results on uniqueness and other properties of ground states within the framework of quantum mechanics. The properties of these isolated bodies are often inspiring and provide technical tools required to handle the two-body case. These properties were hence later synthesized and extended by McCann [46]. However, McCann did not provide detailed proofs, and some of the propositions are not entirely straightforward to justify.
This paper aims to achieve the following two goals:
  1. (a)
    To rigorously justify and extend the classical results. We revisit the existence and structure of non-rotating stars given by Auchmuty and Beals [4] and McCann [46] (Theorem 2.6). Following a suggestion of McCann [46], we prove uniqueness result in the classical Newtonian mechanical setting by adapting the quantum‑mechanical variational framework of Lieb and Yau [42] to this setting (Theorem 2.42).
  2. (b)
    To explore the dependence of properties on the total mass of the system under polytropic law by employing a scaling method (Theorem 3.2 and Proposition 3.4). The scaling analysis not only clarifies relations among solutions of different mass but also provides quantitative convergence results in the small-mass limit. In particular, we examine the decay rates of density profiles and the contraction or extension behavior of the supports of solutions (Remark 3.3).
The results of this paper are adapted from the author’s Master’s thesis [15], which also gave rise to two companion papers: [17] and [16]. The present work holds intrinsic interest within the theory of self-gravitating fluids. Furthermore, serving as a bridge within the thesis, this work builds upon conclusions from [17] and, in turn, provides essential a priori estimates for the study of rotating star–planet systems in [16], where comparisons to the non-rotating case are essential. In [17], the conversion from the Euler–Lagrange equation to the Euler–Poisson equation was rigorously established, with particular attention given to showing that the energy-minimizing solutions still satisfy the Euler–Poisson equations at their boundaries. Nevertheless, this work is largely self-contained. Results from [17] are given with proof outlines herein or cited as isolated technical details with complete proofs in [17].
The paper is organized according to the goals above. In Section 2, we introduce the mathematical settings and accomplish goal (a), reviewing and extending the classical results. In Section 3, we address goal (b) through a scaling analysis.
2 Construction and Results of Non-Rotating Systems
In this section we revisit and adapt the results theorems for non-rotating stars given by Auchmuty and Beals [4] and Lieb and Yau [42], based on McCann’s or the author’s construction and assumptions [46] [17]. In the first subsection, we introduce the problem settings and some basic properties. In the second subsection, we review the existence theorems for non-rotating stars given by Auchmuty and Beals [4] and McCann [46]. In the third subsection, we adapt the uniqueness results from the quantum mechanical framework of Lieb and Yau [42] to the classical Newtonian mechanical setting.
2.1 Problem settings
The state of a fluid may be represented by its mass density $\rho(x)\geq 0$ and velocity vector field $v(x)=0$ on $\mathbb{R}^{3}$ for non-rotating case. The fluid interacts with itself through Newtonian gravity hence we need to consider gravitational interaction energy, which will be given later. Moreover, to define internal energy, we consider a general form of the pressure $P(\rho)$ as the following (which is consistent with [17], see also [4, 46, 33]):
Assumption 2.1.
  1. (F1)
    $P:[0,\infty)\rightarrow[0,\infty)$ continuous and strictly increasing;
  2. (F2)
    $\lim\limits_{s\rightarrow 0}P(s)s^{-\frac{4}{3}}=0$;
  3. (F3)
    $\lim\limits_{s\rightarrow\infty}P(s)s^{-\frac{4}{3}}=\infty$.
Results under a weakened version of (F3), denoted as (F3’), will also be discussed later; see Remark 2.13. Here assumption (F3’) is given as follows:
$\lim\limits_{s\rightarrow\infty}\inf P(s)s^{-\frac{4}{3}}>K$ (F3’)
where $K>0$.
With these assumptions, we also define $A(s)$ as the following:
$A(s):=\int_{1}^{\infty}P\left(\frac{s}{v}\right)dv=s\int_{0}^{s}P(\tau)\tau^{-2}d\tau$ (1)
Assuming that $\sigma$ is an energy minimizer, in order to show $\sigma$ is the solution to (EP’) in the classical sense, we want the differentiability of $\sigma$ or of $P(\sigma)$, hence we introduce the following assumption:
  1. (F4)
    $P(\rho)$ is continuously differentiable on $0,\infty)$, and $P^{\prime}(\rho)>0$ if $\rho>0$.
Actually in [17] we introduce another more general but also somewhat technical assumption:
  1. (F4’)
    $P(\rho)$ is continuously differentiable on $0,\infty)$. If $\rho>0,P(\rho)$ has non-vanishing (first order or higher order) derivative at $\rho$. That is, $\exists n\geq 1$, such that $P^{(n)}(\rho)$ exists and is not 0.
Due to (F2) we know $A(s)$ is well-defined. We also list some properties of $A(s)$ from another paper by the author [17, Section 2].
Lemma 2.2 (Properties of $A$ [17, Section 2]).
Let $A(s)$ is defined above (1), then
  1. (ii)
    $A(s)$ is strictly increasing;
  2. (ii)
    AA also satisfies (F2) and (F3) (or (F3’) if we assume $P$ satisfies (F3’) instead of (F3), though the constant $K$ can be different);
  3. (iii)
    $A^{\prime}$ satisfies: $\lim\limits_{s\rightarrow 0}A^{\prime}(s)\rho^{-\frac{1}{3}}=0$ and $\lim\limits_{s\rightarrow\infty}A^{\prime}(s)s^{-\frac{1}{3}}=\infty$;
  4. (iv)
    $A^{\prime}(s)$ is continuous and
    $A^{\prime}(s)=\left\{\begin{array}[]{r}\int_{0}^{s}P(t)t^{-2}dt+\frac{P(s)}{s},s>0\\ 0,s=0\end{array}\right$. (2)
    Moreover, $A^{\prime}(s)s-A(s)=P(s)$, $A^{\prime\prime}(s)=\frac{P^{\prime}(s)}{s}\geq 0$ a.e.;
  5. (v)
    $A^{\prime}$ is continuous and strictly increasing. Moreover, $A$ is convex;
  6. (vi)
    The inverse function of $A^{\prime}$, denoted by $\left(A^{\prime}\right)^{-1}$ or $\phi$, is well defined on $0,\infty)$. Moreover, $\phi=\left(A^{\prime}\right)^{-1}$ is continuous since $A^{\prime}$ is continuous;
  7. (vii)
    If we assume $P(\rho)$ satisfies (F4), then $A^{\prime\prime}(\rho)$ exists and $A^{\prime\prime}(\rho)=\frac{P^{\prime}(\rho)}{\rho}\neq 0$ is continuous if $\rho>0$. In particular, $\phi=\left(A^{\prime}\right)^{-1}\in C^{1}((0,\infty))$.
If polytropic equations of state $P(s)=Ks^{\gamma}$ holds, where the parameter $\gamma>\frac{4}{3}$, then easy to check $P$ satisfies (F1) (F2) (F3)(F4) or (F1)(F2)(F3)(F4’), and
$A(s)=\frac{K}{\gamma-1}s^{\gamma}$ (3)
Hence one can check $A(s)$ indeed satisfies Lemma 2.2.
In the following sections, we assume $P(\rho)$ satisfies (F1)(F2)(F3) unless otherwise specified. We will mention (F4) or (F4’) or other assumptions when we want to use them.
We consider an “admissible class” for $\rho$ as the following:
$R\left(\mathbb{R}^{3}\right)$ $:=\left\{\left.\rho\in L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)\right\rvert\,\rho\geq 0,\int_{\mathbb{R}^{3}}\rho\,dx=1\right\}$ (4)
Then given $\rho$ in such set, we consider energy $E_{0}(\rho)$ consisting of two terms:
$E_{0}(\rho):=U(\rho)-\frac{G(\rho,\rho)}{2}$ (5)
$U(\rho):=\int_{\mathbb{R}^{3}}A(\rho(x))\,dx$ (6)
$G(\sigma,\rho):=\int_{\mathbb{R}^{3}}V_{\sigma}\rho\,dx=\iint_{{\mathbb{R}^{3}}\times{\mathbb{R}^{3}}}\frac{\rho(x)\sigma(y)}{|x-y|}\,dxdy$ (7)
Here $U(\rho)$ is the internal energy, and $G(\rho,\rho)$ is the gravitational potential energy (also called gravitational interaction energy).
Remark 2.3.
In fact, Auchmuty and Beals [4] consider a more general form of the energy with angular momentum
$E(\rho)=E_{0}(\rho)+\frac{1}{2}\int_{\mathbb{R}^{3}}\rho(x)L\left(m_{\rho}(r(x))\right)r^{-2}(x)\,dx$ (8)
Another model for uniformly rotating stars was studied by McCann [46], where the total angular momentum $\boldsymbol{J}=J\hat{e}_{z}=(0,0,J)^{T}$ is prescribed. The corresponding energy $E_{J}(\rho)$ is defined as the following:
$E_{J}(\rho)=U(\rho)-\frac{G(\rho,\rho)}{2}+T_{J}(\rho)$ (9)
with the kinetic energy $T_{J}(\rho)$ given by
$T_{J}(\rho):=\frac{J^{2}}{2I(\rho)}$ (10)
In this paper we mainly consider the non-rotating case ($L=0$ or $J=0$, reflected in the subscript 0{0} in $E_{0}(\rho)$). This is not only for the sake of simplicity but also helps to produce more results such as uniqueness result in subsection 2.3.
Remark 2.4.
Since $\rho\in L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)\cap L^{1}\left(\mathbb{R}^{3}\right)$, we have $G(\rho,\rho)<\infty$, see for example [17, Section 2]. Hence $E_{0}(\rho)$ is well-defined.
Actually we can generalize our definition of $E_{0}(\rho)$ to the set
$mR\left(\mathbb{R}^{3}\right)$ $=\left\{\left.\rho\in L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)\right\rvert\,\rho\geq 0,\int_{\mathbb{R}^{3}}\rho\,dx=m\right\}$ (11)
Remark 2.5.
From Remark 2.4, we see one reason why we require $\rho$ to belong not only to $L^{1}$ but also to $L^{\frac{4}{3}}$. Furthermore, selecting the specific exponent $\frac{4}{3}$ can also be shown as a natural assumption if we hope the minimal energy $\inf\limits_{\rho\in m{R}\left(\mathbb{R}^{3}\right)}E_{0}(\rho)$ to be finite. We will discuss this later (Proposition 2.16, see also [4, Section 1,6 and 8], and [46, Section 2]). We will also show actually $\frac{4}{3}$ is the critical number such that the maximal energy is finite (Corollary 2.18).
Consider the non-rotating minimizer $\sigma_{m}$ of $E_{0}(\rho)$ among configurations of mass $m\in[0,\infty)$, the corresponding minimum energy is finite due to the remark above. For the sake of convenience, we denote them by
$e_{0}(m):=E_{0}\left(\sigma_{m}\right)=\inf_{\rho\in{R}\left(\mathbb{R}^{3}\right)}E_{0}(m\rho).$ (12)
We also denote $e_{0}(1)$ by $e_{0}$.
Since the energy is translation-invariant, we may just consider looking for an energy minimizer $\rho$ such that the center of mass $\bar{x}(\rho)$ is 0, where
$\bar{x}(\rho):=\frac{\int_{\mathbb{R}^{3}}x\rho(x)\,dx}{\int_{\mathbb{R}^{3}}\rho(x)\,dx}$ (13)
Additionally, let support of $\rho$ be the smallest closed set carrying the full mass of $\rho$ (denoted by spt $\rho$). Intuitively we hope spt $\rho$ to be compact and simply-connected, as this aligns with the case of a single-star system.
2.2 Existence Results for Non-rotating Bodies
Recall that in the calculus of variations, if an energy functional has a minimizer, then this minimizer must satisfy the Euler–Lagrange equation [^1] When the minimization is carried out over a restricted class, the Euler–Lagrange equation may take the form of an inequality or involve taking the positive part; see, for example, [17, Section 2, Appendix A].. Moreover, it turns out that the Euler–Lagrange equation is equivalent, in a certain sense, to the Euler–Poisson equation (see, for example, [17, Section 2] or [33]). Based on this observation, we have Auchmuty, Beals, Lieb and Yau’s results [4, 42]. They not only show the existence of one-body and non-rotating solutions of Euler-Poisson equations, but also indicate some useful properties. McCann lists those results as the following theorem:
Theorem 2.6 (Non-rotating Stars [4, 42, 46]).
For $E_{0}(\rho)$ from (5), $e_{0}(m)$ from (12) and $m\in[0,\infty)$:
  1. (i)
    $E_{0}(\rho)$ attains its minimum $e_{0}(m)$ among $\rho$ such that $\rho\in m{R}\left(\mathbb{R}^{3}\right)$;
  2. (ii)
    $e_{0}(m)$ decreases continuously from $e_{0}(0)=0$ and is strictly concave;
There are bounds $R_{0}(m)$ and $C_{0}(m)$ on the radius and central density, such that any mass $m$ minimizer $\sigma_{m}$ of $E_{0}(\rho)$ satisfies.
  1. (iii)
    $\sigma_{m}$ is spherically symmetric and radially decreasing after translation;
  2. (iv)
    $\left\|\sigma_{m}\right\|_{L^{\infty}}\leq C_{0}(m)$;
  3. (v)
    spt $\sigma_{m}$ is contained in a ball of radius $R_{0}(m)$;
  4. (vi)
    $\sigma_{m}$ is continuous; moreover, if $P(\rho)$ satisfies (F4), then $\phi\in C^{1}((0,\infty))$ and $\sigma_{m}\in C^{1}(\{\sigma_{m}>0\})$;
  5. (vii)
    $\sigma_{m}$ satisfies
    $A^{\prime}(\sigma_{m}(x))=\left[V_{\sigma_{m}}(x)+\lambda_{m}\right]_{+}$ (EL)
    on all of $\mathbb{R}^{3}$ and a single Lagrange multiplier $\lambda_{m}\left\{\begin{array}[]{l}<0,m>0\\ =0,m=0\end{array}\right$.. Here $\cdot$$_{+}$ is the nonnegative (positive) part function defined as $\lambda]_{+}:=\max\{\lambda,0\}$;
  6. (viii)
    the left and right derivatives of $e_{0}(m)$ bound $\lambda_{m}$: $e_{0}^{\prime}\left(m^{+}\right)\coloneq\lim\limits_{\widetilde{m}\rightarrow m^{+}}{e_{0}^{{}^{\prime}}\left(\widetilde{m}\right)}\leq\lambda_{m}\leq e_{0}^{\prime}\left(m^{-}\right)\coloneq\lim\limits_{\widetilde{m}\rightarrow m^{-}}{e_{0}^{{}^{\prime}}\left(\widetilde{m}\right)}$;
  7. (ix)
    if $P(\rho)$ satisfies assumption (F4’), then $\rho$ satisfies the reduced Euler-Poisson equations (EP’).
Remark 2.7.
Theorem 2.6 (iii), which is based on the rearrangement inequality, actually implies the support of $\rho$ is simply connected. This corresponds to a single-star system and thus satisfies our expectation. In contrast, we point out that in McCann’s binary star system [46], while he shows that the support of $\rho$ lies in two separate balls, it does not prove that the support consists of exactly two connected components. Whether it truly has exactly two connected components—and thus corresponds to a binary star system—remains an open question worthy of discussion.
Remark 2.8.
Notice Auchmuty and Beals only concluded that (EP’) is satisfied in the region where $\rho>0$ [4, Theorem A], whereas in Theorem 2.6 (ix) we establish a stronger result—namely, that the equation (EP’) holds over the entire $\mathbb{R}^{3}$ space.
Given that $\rho$ is an energy minimizer, to exploit this property better, we introduce the following perturbation set $P_{\infty}(\rho)$(the motivation behind its definition and related properties can be found in [17, Section 4]), which depends on $\rho$:
$P_{\infty}(\rho):=\bigcup_{R<\infty}P_{R}(\rho)$ (14)
Here, $P_{R}(\rho)$ is defined as:
[…]
One can see $P_{\infty}(\rho)$ is a convex cone. To apply variational method, we need to consider the derivative of $E_{0}$:
Lemma 2.9 (Differentiability of Energy $E_{0}(\rho)$ [17, Section 5]).
Given $\rho\in m{R}^{3}$ with $U(\rho)<\infty$, we have $E_{0}(\rho)$ is differentiable at $\rho$ in the direction of $P_{\infty}(\rho)$. The derivative at $\rho$ is $E_{0}^{\prime}(\rho)$ in the sense that $\forall\sigma\in P_{\infty}(\rho)$, $E_{0}^{\prime}(\rho)(\sigma)=\int_{\mathbb{R}^{3}}E_{0}^{\prime}(\rho)\sigma\,dx$ [^2] To remain consistent with the notation in [4, Section 4], we use $E_{0}^{\prime}$ as the symbol for both linear functional and function, provided it does not cause confusion., where on the right hand side the function $E_{0}^{\prime}(\rho)$ is given by
$E_{0}^{\prime}(\rho)(x):=A^{\prime}(\rho(x))-V_{\rho}(x)$ (15)
Remark 2.10.
To apply Lemma 2.9, we need to verify $U(\rho)<\infty$. It turns out if $\rho\in L^{1}\cap L^{\infty}$, or $\rho$ minimizes $E_{0}(\rho)$ locally (under the topology induced by Wasserstein $L^{\infty}$ distance) or globally on $m{R}\left(\mathbb{R}^{3}\right)$, then $U(\rho)<\infty$, see [17, Section 5].
Proof of Theorem 2.6.
Here we provide a proof outline, supplement some ideas and details not mentioned explicitly in [4, 42]. One can refer to [4, 42] and references therein to understand some step(s) omitted here.
When $m=0$, we can see $\rho=0$, hence the results (i) (iii-vii) are easy to check. The results (ii) and (viii) can be proved in the same way as in the case $m>0$. Therefore, we assume $m>0$ in the following arguments unless otherwise specified.
To prove the existence of a minimizer  [^3] It is natural to consider direct method [19]. However, so far, we only know $mR\left(\mathbb{R}^{3}\right)$ is bounded in $L^{1}\left(\mathbb{R}^{3}\right)$, which only implies weak star convergence. It is not enough to show the weak continuity of $G(\rho,\rho)$ and then the weak lower semi-continuity of $E_{0}(\rho)$ with respect to $\rho$ on $W_{\infty}\coloneq m{R}\left(\mathbb{R}^{3}\right)$, see Proposition A.5 or Proposition A.6. Hence, the direct method may not be valid on $W_{\infty}$. We need to construct a constrained admissible class $W_{R}$., we first construct a constrained admissible class:
[…]
Here $B_{R}(\cdot)$ is the open ball defined by $B_{R}(\cdot):=\left\{x\in\mathbb{R}^{3}\mid|x-\cdot|<R\right\}$. The axial symmetry and centered condition in (16) ensure that the center of mass of $\rho_{R}$ is $(0,0,0)^{T}$. Notice that $B_{R}(0)$ is a bounded set, by Proposition A.5 or Proposition A.6, we can show the weak continuity of $G(\rho,\rho)$ with respect to $\rho$ on $W_{R}$. Then one can apply direct method to show the existence of a minimizer $\sigma_{R}$ (depending on $m$) on $W_{R}$. In order to show such minimizer $\sigma_{R}$ is truly a minimizer on the whole admissible class $W_{\infty}$ for sufficiently large $R$, we need the following 4 lemmas.
  1. (1)
    Show the $L^{\infty}\left(\mathbb{R}^{3}\right)$ bound of $\sigma_{R}$ is uniform, i.e. $\forall R>R_{0},\left\|\sigma_{R}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\leq k_{1}$. (Lemma 2.21)
  2. (2)
    Show $\forall R>\widetilde{R}_{0}$, $E_{0}\left(\sigma_{R}\right)\leq\alpha<0$, $\lambda_{R}\leq l<0$, where $\lambda_{R}$ is the corresponding Lagrange multiplier. (Lemma 2.24)
  3. (3)
    Show $\forall R>R_{1}$, spt $\sigma_{R}$ is contained in $\overline{B_{R_{1}}(0)}$. (Lemma 2.27)
  4. (4)
    Take $R^{\prime}=\max\left\{k_{1},R_{1}\right\}+1$, show $\sigma_{R^{\prime}}$ is a minimizer on $W_{\infty}=m{R}\left(\mathbb{R}^{3}\right)$. (Lemma 2.28)
The proofs of the above four lemmas are postponed to later text. Once these lemmas are established, we obtain a minimizer which satisfies Parts (iv) (v). Therefore, Part (i) holds true with finite $e_{0}(m)$ by the second lemma. With the strategies similar to that employed in the proofs of the above lemmas, one can show Part (iv) and Part (v) actually hold true for any minimizer $\sigma_{m}$ (observe that the last term in (25) can be replaced by $\frac{m}{4R}$ since $\int_{B_{R}(0)}\sigma_{m}\,dx\geq\frac{m}{2}$ for sufficiently large $R$). More precisely, the proof of Part (v) also relies on Part (vi) to show that outside the ball of radius $R_{0}(m)$, $\sigma_{m}$ is not only almost everywhere zero but truly zero everywhere. Moreover, we have the fact that the Lagrange multiplier $\lambda_{m}$ is negative, which follows from the similar arguments in the proof of Step (1) (Lemma 2.21) showing $\lambda_{R}<0$, see also Remark 2.25.
Now we have the existence of the minimizer $\sigma_{m}$. Since $\sigma_{m}$ is a minimizer on $m{R}\left(\mathbb{R}^{3}\right)$, the set of admissible perturbations here is exactly $P_{\infty}\left(\sigma_{m}\right)$. One can apply (generalized) Lagrange multiplier theorem ([4, Section 2] or [17, Appendix A]), together with the fact that there are sufficiently many functions in $P_{\infty}(\sigma_{m})$, to obtain the following relations (see also [4, Lemma 2] or [17, Section 5]):
$E_{0}^{\prime}\left(\sigma_{m}\right)\geq\lambda_{m}\text{ a.e., }$ (17)
and
$E_{0}^{\prime}\left(\sigma_{m}\right)=\lambda_{m}\text{ a.e. on }\{\sigma_{m}(x)>0\}.$ (18)
Since $E_{0}^{\prime}\left(\sigma_{m}\right)(x)=A^{\prime}\left(\sigma_{m}(x)\right)-V_{\sigma_{m}}(x)$, (18) implies
$A^{\prime}\left(\sigma_{m}(x)\right)\leq V_{\sigma_{m}}(x)+\lambda_{m}\text{ a.e. on }\{\sigma_{m}(x)>0\}$ (19)
Due to Lemma 2.2 (iii), notice $\lambda_{m}<0$, we can obtain $\sigma_{m}{}^{\frac{1}{3}}\leq cV_{\sigma_{m}}$ a.e. on the set where $\sigma_{m}\geq K$ for some $K$ large enough. As in [17, Section 2], we can show the bound of $\|\sigma_{m}\|_{L^{p}}$ and $\left\|V_{\sigma_{m}}\right\|_{L^{\widetilde{p}}}$ with $p$ and $\widetilde{p}$ increasing alternately and reaching infinity in finite steps, and then $V_{\sigma_{m}}$ is continuously differentiable (bootstrap method). On the other hand (17)(18) implies
$\sigma_{m}=\phi\circ\left[V_{\sigma_{m}}(x)+\lambda_{m}\right]_{+}\text{ a.e. in }\mathbb{R}^{3}$ (20)
where $\phi=\left(A^{\prime}\right)^{-1}$ is continuous. This means $\sigma_{m}$ has a continuous representative. Moreover, if $P(\rho)$ satisfies (F4), then $\phi\in C^{1}((0,\infty))$ by Lemma 2.2 (vii), which implies $\sigma_{m}\in C^{1}\left(\left\{\sigma_{m}>0\right\}\right)$ as shown in [17, Section 2]. (Part (vi)).
Part (iii) comes from a strong rearrangement inequality in Lieb [43, Lemma 3] and the generalization by Sobolev of the Hardy-Littlewood theorem on rearrangements of functions [54]. One can also refer to [46, Proof of Theorem 6.1] or [16, Section 6]. (Part (iii))
For Part (vii), we already know $\lambda_{m}<0$. Moreover, (17) (18), together with $\sigma_{m}$ is continuous, give (EL), see for example [46, Section 5] or [17, Section 2]. (Part (vii))
For Part (ii), we only consider the case that complies with polytropic equations of state $P(\sigma)=K\sigma^{\gamma}$. The general case can be found in [42, Theorem 3 (d)] [^4] However, it should be noted that the argument there is set within a quantum mechanical framework and does not include an internal energy term, although McCann mentioned that the results can be applied “equally well to all $A(\rho)$ consistent with (F1)-(F3)[46, Section 3].. We utilize the scaling relations mentioned in Section 3. That is, given $\sigma$ a non-rotating minimizer with mass $1$, then $\sigma_{m}(x)=\frac{1}{A}\sigma\left(\frac{1}{B}x\right)$ will be a non-rotating minimizer with mass $m$, where $A(m)=m^{-\frac{2}{3\gamma-4}}$, $B(m)=m^{\frac{\gamma-2}{3\gamma-4}}$. At the same time, we know $e_{0}(m)=E_{0}\left(\sigma_{m}\right)=E_{0}\left(\frac{1}{A}\sigma\left(\frac{1}{B}x\right)\right)=m^{\frac{5\gamma-6}{3\gamma-4}}e_{0}$. Notice when $m=0$, such expression also holds since $e_{0}(0)=0$. By the similar arguments as the proof of Step (2) (Lemma 2.24), we know $e_{0}\coloneq e_{0}(1)<0$, which implies $e_{0}(m)$ decreases continuously from $e_{0}(0)=0$ and is strictly concave (Part (ii)).
For Part (viii), first the concavity of $e_{0}(m)$ implies the existence of left-hand and right-hand derivations (since they are monotonic), and they are equal a.e. (since concavity implies Lipschitz continuity and then apply Rademacher’s theorem [21, Subsection 5.8.3]). Thanks to Euler-Lagrange equations (EL), we multiply $\sigma_{m}$ on both sides and integrate, and then we obtain
$\int_{\mathbb{R}^{3}}E_{0}^{\prime}\left(\sigma_{m}\right)\sigma_{m}\,dx=\lambda_{m}\cdot m$
Notice due to Part (iv) and (v), we know $\sigma_{m}\in P_{\infty}\left(\sigma_{m}\right)$ thus $\int_{\mathbb{R}^{3}}E_{0}^{\prime}\left(\sigma_{m}\right)\sigma_{m}\,dx$ makes sense and
$\int_{\mathbb{R}^{3}}E_{0}^{\prime}\left(\sigma_{m}\right)\sigma_{m}\,dx=\lim\limits_{t\rightarrow 0^{+}}\frac{E_{0}\left(\sigma_{m}+t\sigma_{m}\right)-E_{0}\left(\sigma_{m}\right)}{t}=\lim\limits_{t\rightarrow 0^{-}}\frac{E_{0}\left(\sigma_{m}\right)-E_{0}\left(\sigma_{m}-t\sigma_{m}\right)}{t}$
Since $\sigma_{m}+t\sigma_{m}$ has mass $(1+t)m$, then
$e_{0}((1+t)m)=E_{0}\left(\sigma_{(1+t)m}\right)\leq E_{0}\left(\sigma_{m}+t\sigma_{m}\right)$
Thus
$\lim\limits_{t\rightarrow 0^{+}}\frac{E_{0}\left(\sigma_{m}+t\sigma_{m}\right)-E_{0}\left(\sigma_{m}\right)}{t}\geq\lim\limits_{t\rightarrow 0^{+}}\frac{e_{0}((1+t)m)-e_{0}(m)}{t}=me_{0}^{\prime}\left(m^{+}\right)$
Hence we have $\lambda_{m}\geq e_{0}^{\prime}\left(m^{+}\right)$. Similarly, we have $\lambda_{m}\leq e_{0}^{\prime}\left(m^{-}\right)$. (Part (viii))
For Part (ix), thanks to Part (vi), we know the global continuity of $\sigma_{m}$, and then we can apply the same arguments as in [17, Section 2] and show the existence of $\nabla P\left(\sigma_{m}\right)$ on the boundary $\partial\left\{\sigma_{m}>0\right\}$, and verify $\sigma_{m}$ is a solution to reduced Euler-Poisson equations (EP’) in $\mathbb{R}^{3}$. (Part (ix)) ∎
Remark 2.11.
In the proof of Part (iv) of Theorem 2.6, the relaxation of the last term in (25) to $\frac{m}{4R}$ is not fully justified in [4]; here we make this detail explicit for completeness.
Remark 2.12.
When we prove Theorem 2.6 (vi), we mention $\sigma_{m}{}^{\frac{1}{3}}\leq cV_{\sigma_{m}}$ a.e. on the set where $\sigma_{m}\geq K$ for some $K$ large enough. Actually it turns out even if we do not know $\lambda_{m}<0$, $\sigma_{m}{}^{\frac{1}{3}}\leq cV_{\sigma_{m}}$ can still hold true since we can absorb $\lambda_{m}$ when $\sigma_{m}$ is large enough. See [4, Lemma 3].
Remark 2.13.
We point out in Theorem 2.6 we can show the existence of non-rotating single star no matter how large the mass is. However, if we replace (F3) by (F3’), then we can only obtain minimizers with total mass not larger than $m(K)$ via variational method for some $m(K)>0$. $m(K)$ here turns out to be the Chandrasekhar mass for the model. If the mass is too large, gravitational collapse may happen [4, Appendix].
To clarify, replacing (F3) with (F3’) would require more careful treatment to prove Lemma 2.6. First, observe that Lemma 2.2 needs to be modified under assumption (F3’) rather than (F3). In the proof of Lemma 2.6, a potential issue with a large mass is that it may invalidate the earlier assertion that $\sigma_{m}{}^{\frac{1}{3}}\leq cV_{\sigma_{m}}$ a.e. on the set where $\sigma_{m}\geq K$ for some $K$ large enough”. Moreover, (F3’) may also lead to a more delicate treatment in the proof of Lemma 2.21 below; see also Remark 2.23. For more discussion, one can see [46, Section 2] or [4, Section 1,6,8 and Appendix].
Remark 2.14.
Theorem 2.6 (ii) actually implies that $e^{\prime}_{0}(m^{-})<0$ for $m>0$. To prove it, first note $e_{0}(m)<0$ for $m>0$. Otherwise the monotonicity would imply that $e_{0}=0$ on $[0,m]$, which contradicts the strict concavity property. Then due to strict concavity, $\forall m>0$, $\forall\epsilon>0$, $\frac{e(m+\epsilon)-e(m)}{\epsilon}<\frac{e(m)-e(0)}{m}=\frac{e(m)}{m}<0$, which implies $e^{\prime}_{0}(m^{-})\leq\frac{e(m)}{m}<0$.
Remark 2.15.
Currently, we do not know whether the minimizer $\sigma_{m}$ with fixed mass $m$ is unique. Further discussion regarding uniqueness (up to translation) will be conducted under additional assumptions later. However, even if multiple minimizers $\sigma_{m}^{1},\sigma_{m}^{2}\ldots\sigma_{m}^{i}\ldots$ exist under the present assumptions, due to Theorem 2.6 (iv), together with interpolation inequality [8, Section 4.2], we know that all of them have uniform $L^{p}$ bound. Moreover, according to Theorem 2.6 (viii) and Remark 2.14, we observe their corresponding Lagrange multipliers $\lambda_{m}^{i}$ are also uniformly bounded by $e^{\prime}_{0}(m^{-})$, which is strictly negative. This observation is applied in [46, Section 6].
Before showing the existence of a minimizer on $m{R}\left(\mathbb{R}^{3}\right)$, it is worth noting that if we replace $\frac{4}{3}$ by any smaller number in (F3), there is no minimizer since the energy may be unbounded below. The idea comes from [4, Remark in Section 8].
Proposition 2.16 (Unbounded below if $\gamma$ $\frac{4}{3}$).
If $P(\rho)$ satisfies $\lim\limits_{\rho\rightarrow\infty}\sup P(\rho)\rho^{-\gamma}=K$ instead of (F3), where $\gamma<\frac{4}{3},0\leq K<\infty$, then the energy $E_{0}(\rho)$ is unbounded below, that is $\inf\limits_{\rho\in m{R}\left(\mathbb{R}^{3}\right)}E_{0}(\rho)=$-$\infty$. In particular, there is no minimizer on $m{R}\left(\mathbb{R}^{3}\right)$.
Proof.
Follow the proof of L’Hôpital’s rule [52, Theorem 5.13], we know $\lim\limits_{\rho\rightarrow\infty}\sup P(\rho)\rho^{-\gamma}=K$ implies $\lim\limits_{\rho\rightarrow\infty}\sup A(\rho)\rho^{-\gamma}\leq 3K$. then we know $\lim\limits_{\rho\rightarrow\infty}A(\rho)\rho^{-\frac{4}{3}}=$ 0. Therefore, $\forall\hat{\epsilon}>0,\exists S_{\hat{\epsilon}}>0$, such that $A(s)\leq\hat{\epsilon}s^{-\frac{4}{3}}$ when $s\geq S_{\hat{\epsilon}}$. In the statement of Proposition A.2, let $\hat{c}$ be the smallest constant such that $\forall\rho\in L^{1}\cap L^{\frac{4}{3}}$, we have
$\left|\int_{\mathbb{R}^{3}}\rho V_{\rho}\,dx\right|\leq 2\hat{c}\left(\int_{\mathbb{R}^{3}}|\rho|\,dx\right)^{\frac{2}{3}}\int_{\mathbb{R}^{3}}|\rho|^{\frac{4}{3}}\,dx$
Then $\hat{c}>0$. Otherwise $\left|\int_{\mathbb{R}^{3}}\rho V_{\rho}\,dx\right|\leq 0$ implies $\rho=0$, but 0 is not the only element in $L^{1}\cap L^{\frac{4}{3}}$.
[…]
$\left|\int_{\mathbb{R}^{3}}\widetilde{\rho_{\epsilon}}V_{\widetilde{\rho_{\epsilon}}}\,dx\right|>2\left(\hat{c}-\frac{\epsilon}{2}\right)\left(\int_{\mathbb{R}^{3}}\left|\widetilde{\rho_{\epsilon}}\right|\,dx\right)^{\frac{2}{3}}\int_{\mathbb{R}^{3}}\left|\widetilde{\rho_{\epsilon}}\right|^{\frac{4}{3}}\,dx$
[…]
$\frac{\left|\int_{\mathbb{R}^{3}}\rho_{\epsilon}V_{\rho_{\epsilon}}\,dx\right|}{\left(\int_{\mathbb{R}^{3}}\left|\rho_{\epsilon}\right|\,dx\right)^{\frac{2}{3}}\int_{\mathbb{R}^{3}}\left|\rho_{\epsilon}\right|^{\frac{4}{3}}\,dx}>2(\hat{c}-\epsilon)$
We can also write it as
$G\left(\rho_{\epsilon},\rho_{\epsilon}\right)=\left|\int_{\mathbb{R}^{3}}\rho_{\epsilon}V_{\rho_{\epsilon}}\,dx\right|>2(\hat{c}-\epsilon)\left(\int_{\mathbb{R}^{3}}\left|\rho_{\epsilon}\right|\,dx\right)^{\frac{2}{3}}\int_{\mathbb{R}^{3}}\left|\rho_{\epsilon}\right|^{\frac{4}{3}}\,dx$
It implies $\rho_{\epsilon}\neq 0$. We can also assume $\rho_{\epsilon}\geq 0$ since we can replace $\rho_{\epsilon}$ by $\left[\rho_{\epsilon}\right]_{+}$ if needed. Without loss of generality, we assume $\int_{\mathbb{R}^{3}}\rho_{\epsilon}\,dx=1$. Let $\rho_{\delta}(x)=m\delta^{-3}\rho_{\epsilon}\left(\delta^{-1}x\right)$, $\delta>0$. Then $\rho_{\delta}\in m{R}\left(\mathbb{R}^{3}\right)$ since $C_{c}^{0}\left(\mathbb{R}^{3}\right)\subset L^{\frac{4}{3}}$. Similar as discussions in Section 3, we can show $G\left(\rho_{\delta},\rho_{\delta}\right)=\frac{m^{2}}{\delta}G\left(\rho_{\epsilon},\rho_{\epsilon}\right)$, and
$U\left(\rho_{\delta}\right)$ $=\int_{\mathbb{R}^{3}}A\left(\rho_{\delta}\right)\,dx$
$=\int_{\left\{\rho_{\delta}\geq S_{\hat{\epsilon}}\right\}}A\left(\rho_{\delta}\right)\,dx+\int_{\left\{\rho_{\delta}<S_{\hat{\epsilon}\}}\right.}A\left(\rho_{\delta}\right)\,dx$
$\leq\int_{\left\{\rho_{\delta}\geq S_{\hat{\epsilon}}\right\}}\hat{\epsilon}\cdot\rho_{\delta}^{\frac{4}{3}}\,dx+\int_{\left\{\rho_{\delta}<S_{\hat{\epsilon}\}}\right.}A\left(S_{\hat{\epsilon}}\right)\,dx$
[…]
Here $\mu$ denotes the Lebesgue measure. Notice that given $\hat{\epsilon}$ and $\epsilon$, as $\delta\rightarrow 0$, we have
$\mu\left(\left\{\rho_{\delta}<S_{\hat{\epsilon}}\right\}\right)=\mu\left(\left\{x\left\lvert\,\rho_{\epsilon}\left(\delta^{-1}x\right)<\frac{S_{\hat{\epsilon}}\delta^{3}}{m}\right.\right\}\right)=\delta^{3}\mu\left(y\left\lvert\,\left\{\rho_{\epsilon}(y)<\frac{S_{\hat{\epsilon}}\delta^{3}}{m}\right\}\right.\right)\rightarrow 0$
Then we pick $\epsilon=\frac{\hat{c}}{2},\hat{\epsilon}=\frac{m^{\frac{2}{3}}\hat{c}}{4}$, and then when $\delta\rightarrow 0$
$E_{0}\left(\rho_{\delta}\right)$ $=U\left(\rho_{\delta}\right)-\frac{G\left(\rho_{\delta},\rho_{\delta}\right)}{2}$
$\leq\frac{m^{2}\hat{c}}{4\delta}\int_{\mathbb{R}^{3}}\rho_{\epsilon}^{\frac{4}{3}}\,dx+A\left(S_{\hat{\epsilon}}\right)\mu\left(\left\{\rho_{\delta}<S_{\hat{\epsilon}}\right\}\right)-\frac{m^{2}\hat{c}}{2\delta}\int_{\mathbb{R}^{3}}\rho_{\epsilon}^{\frac{4}{3}}\,dx\rightarrow$-$\infty$
Therefore, $\inf\limits_{\rho\in m{R}\left(\mathbb{R}^{3}\right)}E_{0}(\rho)=$-$\infty$. ∎
Remark 2.17.
Actually, there is another way to see the unboundedness. From $\lim\limits_{\rho\rightarrow\infty}\sup A(\rho)\rho^{-\gamma}\leq 3K$, we can notice $A(\rho)\leq(3K+\hat{\epsilon})\rho^{\gamma}$ for large $\rho$, then similarly as above we have
$E_{0}\left(\rho_{\delta}\right)$ $=U\left(\rho_{\delta}\right)-\frac{G\left(\rho_{\delta},\rho_{\delta}\right)}{2}$
$\leq m^{\gamma}\delta^{-3+3\gamma}(3K+\hat{\epsilon})\int_{\mathbb{R}^{3}}\rho_{\epsilon}^{\gamma}\,dx+A\left(S_{\hat{\epsilon}}\right)\mu\left(\left\{\rho_{\delta}<S_{\hat{\epsilon}}\right\}\right)-\frac{(\hat{c}-\epsilon)m^{2}}{\delta}\int_{\mathbb{R}^{3}}\left|\rho_{\epsilon}\right|^{\frac{4}{3}}\,dx$
Notice $3+3\gamma<-1$, thus we can choose suitable $\epsilon$ and $\hat{\epsilon}$ such that $E_{0}\left(\rho_{\delta}\right)\rightarrow$-$\infty$ as $\delta\rightarrow 0$. From this point we can know that the number $\gamma=\frac{4}{3}$ in (F3) is chosen such that $3+3\gamma\geq-1$, then (F3) can help the blowing up of $U(\rho)$ to control the gravitational collapse $G(\rho,\rho)$.
In fact, in the case that blowing up of $U(\rho)$ is too large, with the same strategy we can show the energy may also be unbounded above. That is, if $\frac{4}{3}$ is replaced by any larger number in (F3), then we will have $U\left(\rho_{\delta}\right)\gg G\left(\rho_{\delta},\rho_{\delta}\right)\gg 1$.
Corollary 2.18 (Unbounded above if $\gamma>\frac{4}{3}$).
If $P(\rho)$ satisfies $\lim\limits_{\rho\rightarrow\infty}\inf P(\rho)\rho^{-\gamma}=K$ instead of (F3), where $\gamma>\frac{4}{3}$, $0<K\leq\infty$, then the energy $E_{0}(\rho)$ is unbounded above, that is $\sup\limits_{\rho\in m{R}\left(\mathbb{R}^{3}\right)}E_{0}(\rho)=+\infty$.
Proof.
Thanks to the arguments above, we can first find a $\rho\in C_{c}^{0}\left(\mathbb{R}^{3}\right)$ with mass 1, and set $\rho_{\delta}(x)=m\delta^{-3}\rho\left(\delta^{-1}x\right)$. If $K<\infty$, notice
$\left\{x\,\middle|\,\rho_{\delta}\geq S_{\hat{\epsilon}}\right\}=\left\{x\,\middle|\,\rho\left(\delta^{-1}x\right)\geq\frac{S_{\hat{e}}\delta^{3}}{m}\right\}$
by a change of variable, we know for $\delta$ large enough, we have
$U\left(\rho_{\delta}\right)\geq m^{\gamma}\delta^{-3+3\gamma}(3K-\hat{\epsilon})\int_{\left\{\rho\geq\frac{S_{\hat{\epsilon}}\delta^{3}}{m}\right\}}\rho^{\gamma}\,dx\geq m^{\gamma}\delta^{-3+3\gamma}(3K-\hat{\epsilon})\int_{\{\rho\geq C\}}\rho^{\gamma}\,dx$
Notice we can find a $C$ such that $\int_{\{\rho\geq C\}}\rho^{\gamma}\,dx>0$ due to the same arguments in [17, Section 4]. By Proposition A.2, together with $\gamma>\frac{4}{3}$, we know as $\delta\rightarrow 0$,
$E_{0}\left(\rho_{\delta}\right)=U\left(\rho_{\delta}\right)-\frac{G\left(\rho_{\delta},\rho_{\delta}\right)}{2}\geq m^{\gamma}\delta^{-3+3\gamma}(3K-\hat{\epsilon})\int_{\{\rho\geq C\}}\rho^{\gamma}\,dx-\frac{\hat{c}m^{2}}{\delta}\int_{\mathbb{R}^{3}}|\rho|^{\frac{4}{3}}\,dx\rightarrow+\infty$
[…]
Remark 2.19.
In fact, the energy may be also unbounded below or above even if we replace ${R}\left(\mathbb{R}^{3}\right)$ by ${R}_{p}\left(\mathbb{R}^{3}\right):=\left\{\rho\in L^{p}\left(\mathbb{R}^{3}\right)\mid\rho\geq 0,\int_{\mathbb{R}^{3}}\rho\,dx=1\right\}$ for some $p\geq 1$, since $C_{c}^{0}\left(\mathbb{R}^{3}\right)\subset L^{p}$ for any $p\geq 1$, and then $\rho_{\epsilon}\in{R}_{p}\left(\mathbb{R}^{3}\right)$.
Remark 2.20.
If (F3) is replaced by the assumption $\lim\limits_{\rho\rightarrow\infty}\inf P(\rho)\rho^{-\frac{4}{3}}=K>0$, Auchmuty and Beals [4, Section 1 and 6] show there is a constant $M_{0}>0$ such that $\forall 0\leq m<M_{0}$ there is a minimizer on $m{R}\left(\mathbb{R}^{3}\right)$. If $\lim\limits_{\rho\rightarrow\infty}\inf P(\rho)\rho^{-\frac{4}{3}}=K>0$ is strengthened to $\lim\limits_{\rho\rightarrow\infty}P(\rho)\rho^{-\frac{4}{3}}=K>0$, then $\forall m>M_{0}$, the energy is unbounded below on $m{R}\left(\mathbb{R}^{3}\right)$ [4, Remark in Section 8]. Moreover, they [4, Remark in Section 6, and Appendix] also show that if $P$ is the function occurring in the theory of white dwarf stars (relativistically degenerate gases), $M_{0}$ turns out to be the same as the limiting mass $M_{3}$ in Chandrasekhar’s theory [12].
We now prove the four steps (lemmas) that are mentioned during the earlier proof of Theorem 2.6. We assume $m>0$, and just provide proof outlines, explain some ideas and add some details not mentioned explicitly in [4, 42]. One can check the proofs in [4] [42] and references therein, to understand some step(s) omitted in the following.
Lemma 2.21 (Uniform bound of $\sigma_{R}$ in $L^{\infty}\left(\mathbb{R}^{3}\right)$ [4]).
Assume $\sigma_{R}$ is the energy minimizer in $W_{R}$, which is given in (16). Then the $L^{\infty}\left(\mathbb{R}^{3}\right)$ bound of $\sigma_{R}$ is uniform, i.e., $\exists C>0$, such that $\forall R>0$, $\left\|\sigma_{R}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\leq C$.
Remark 2.22.
Here we assume that the inequality $\left\|\sigma_{R}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\leq C$ holds automatically if $W_{R}$ is empty. But we will also show $W_{R}$ is not empty when $R$ is large enough in the following proof.
Proof of Lemma 2.21.
[…]
$\int_{\mathbb{R}^{3}}A\left(\sigma_{R}\right)\,dx\leq E_{0}(\widetilde{\rho})+Cm^{\frac{2}{3}}\int_{\mathbb{R}^{3}}\sigma_{R}^{\frac{4}{3}}\,dx$
where $m=\int_{\mathbb{R}^{3}}\sigma_{R}\,dx$. Let $\widetilde{C}=Cm^{\frac{2}{3}}$, by Lemma 2.2 (ii), $\exists s_{0}>0$, such that $\forall s\geq s_{0}$, $(\widetilde{C}+2)s^{\frac{4}{3}}\leq A(s)$. Notice the mass of $\sigma_{R}$ is $m$. Then
$\int_{\mathbb{R}^{3}}\left(\sigma_{R}\right)^{\frac{4}{3}}\,dx$ $=\int_{\left\{x\mid\sigma_{R}(x)<s_{0}\right\}}\left(\sigma_{R}\right)^{\frac{1}{3}}\cdot\sigma_{R}\,dx+\int_{\left\{x\mid\sigma_{R}(x)\geq s_{0}\right\}}\left(\sigma_{R}\right)^{\frac{4}{3}}\,dx$
$\leq s_{0}^{\frac{1}{3}}\cdot m+\frac{1}{(\widetilde{C}+2)}\int_{\mathbb{R}^{3}}A\left(\sigma_{R}\right)\,dx$
[…]
$(\widetilde{C}+1)\int_{\mathbb{R}^{3}}\left(\sigma_{R}\right)^{\frac{4}{3}}\,dx\leq(\widetilde{C}+2)\int_{\mathbb{R}^{3}}\left(\sigma_{R}\right)^{\frac{4}{3}}\,dx-(\widetilde{C}+2)s_{0}^{\frac{1}{3}}\cdot m\leq\int_{\mathbb{R}^{3}}A\left(\sigma_{R}\right)\,dx\leq E_{0}(\widetilde{\rho})+\widetilde{C}\int_{\mathbb{R}^{3}}\left(\sigma_{R}\right)^{\frac{4}{3}}\,dx$
[…]
We claim $\sigma_{R}$ is continuous in $B_{R}(0)$ and the following inequalities hold true (notice $\sigma_{R}(x)>0$ implies $|x|<R$):
$E_{0}{}^{\prime}\left(\sigma_{R}\right)\geq\lambda_{R}\text{ where }|x|<R\text{ and }\sigma_{R}(x)<R$ (21)
$E_{0}{}^{\prime}\left(\sigma_{R}\right)\leq\lambda_{R}\text{ where }\sigma_{R}(x)>0$ (22)
Recall we have $E_{0}^{\prime}\left(\sigma\right)(x)=A^{\prime}\left(\sigma(x)\right)-V_{\sigma}(x)$ by Lemma 2.9, then we can also write
$A^{\prime}\left(\sigma_{R}(x)\right)\leq V_{\sigma_{R}}(x)+\lambda_{R}\text{ where }\sigma_{R}(x)>0$ (23)
In fact, we first apply (generalized) Lagrange multiplier theorem [17, Appendix A] to obtain (21)(22)(23) holds true a.e. in the region described above, then we get
$A^{{}^{\prime}}\left({\sigma_{R}(x)}\right)={\min\left\{{A^{{}^{\prime}}(R),{\max\left\{{V_{\sigma_{R}}(x)+\lambda_{R},0}\right\}}}\right\}}\text{ a.e. in }B_{R}(0)$ (24)
Although the sign of $\lambda_{R}$ is currently unknown (it will be shown negative later), we can still show $\sigma_{R}{}^{\frac{1}{3}}\leq cV_{\sigma_{R}}$ a.e. on the set where $\sigma_{R}\geq K_{R}$ for some $K_{R}>0$ by absorbing $\lambda_{R}$ when $\sigma_{R}$ is large enough [4, Lemma 3]. Similar to the arguments in [17, Section 2], we can show $\sigma_{R}$ is continuous in $B_{R}(0)$ and thereby replace almost everywhere (a.e.) by everywhere.
One may consider using arguments similar to those above (bootstrap method) to show Lemma 2.21. But we need to realize that this time $\sigma_{R}^{\frac{1}{3}}\leq cV_{\sigma_{R}}$ a.e. on the set where $\sigma_{R}\geq K_{R}$, here $K_{R}$ depends on $\lambda_{R}$ thus may not be a uniform constant.
However, this problem can be solved. We claim $K_{R}$ is indeed a uniform constant for sufficiently large $R$. We can show $\lambda_{R}<0$ for sufficiently large $R$. To show $\lambda_{R}<0$, we choose a suitable region $U_{R}$ where (21) holds that $\lambda_{R}\leq E_{0}{}^{\prime}\left(\sigma_{R}\right)$, analyze the behaviors of terms in $E_{0}{}^{\prime}\left(\sigma_{R}\right)$ as $R\rightarrow\infty$, and obtain $\lambda_{R}\leq o\left(R^{-1}\right)-cmR^{-1}<0$ on $U_{R}$. For example, if $x\in B_{R}(0)$, since $\sigma_{R}=0$ outside $B_{R}(0)$, we have
$V_{\sigma_{R}}(x)=\int_{\mathbb{R}^{3}}\frac{\sigma_{R}(y)}{|x-y|}\,dy=\int_{B_{R}(0)}\frac{\sigma_{R}(y)}{|x-y|}\,dy\geq\frac{1}{2R}\int_{B_{R}(0)}\sigma_{R}(y)\,dy=\frac{m}{2R}$ (25)
which gives us the last term $cmR^{-1}$. One can also see how to get the term $o(R^{-1})$ in [4, Lemma 5]. After showing $\lambda_{R}<0$, thanks to (23) we have $A^{\prime}\left(\sigma_{R}\right)\leq V_{\sigma_{R}}$ for $R$ large enough. Recall Lemma 2.2 (iii) tells us $\sigma_{R}^{\frac{1}{3}}\leq A^{\prime}(\sigma_{R})$ when $\sigma_{R}>N$ for some $N>0$. It implies we can take $K_{R}=N$ such that it is a uniform constant for $R\geq R_{1}$, where $R_{1}$ is sufficiently large. Therefore, we can truly apply bootstrap method to prove $\exists C_{1}>0$, such that $\left\|\sigma_{R}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\leq C_{1}$ for $R\geq R_{1}$. Take $C=\max\{C_{1},R_{1}\}$, we know $\left\|\sigma_{R}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\leq C$ for $R>0$. This finishes the proof of Lemma 2.21 (Step (1)). ∎
Remark 2.23.
(F3) is important to show the uniform bound of $\sigma_{R}$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$. It helps to show $\int_{\mathbb{R}^{3}}\left(\sigma_{R}\right)^{\frac{4}{3}}\,dx\leq s_{0}^{\frac{1}{3}}\cdot m+\frac{1}{(\widetilde{C}+2)}\int_{\mathbb{R}^{3}}A\left(\sigma_{R}\right)\,dx$. If we replace $\frac{4}{3}$ by any smaller number in (F3), we could not obtain such uniform bound since we would not have $(\widetilde{C}+1)\int_{\mathbb{R}^{3}}\left(\sigma_{R}\right)^{\frac{4}{3}}\,dx\leq E_{0}(\widetilde{\rho})+\widetilde{C}\int_{\mathbb{R}^{3}}\left(\sigma_{R}\right)^{\frac{4}{3}}\,dx$ as above, as Proposition 2.16 suggests.
To show spt $\sigma_{R}$ is contained in $\overline{B_{R_{1}}(0)}$ (Step (3)), thanks to (23), it’s sufficient to show $V_{\sigma_{R}}(x)\leq-\lambda_{R}$ outside the ball. To do so, we strengthen the result $\lambda_{R}<0$ and obtain $\lambda_{R}\leq l<0$ for all large $R$. We first estimate the energy $E_{0}\left(\sigma_{R}\right)$, which turns out to satisfies $E_{0}\left(\sigma_{R}\right)\leq a<0$ as well.
Lemma 2.24 (Uniform Bound for Energies and Lagrange Multipliers [4]).
$\exists\widetilde{R}_{0},\alpha,l$, such that $\forall R>\widetilde{R}_{0}$, $E_{0}\left(\sigma_{R}\right)\leq\alpha<0$ and $\lambda_{R}\leq l<0$, where $\sigma_{R}$ is the energy minimizer in $W_{R}$, and $\lambda_{R}$ is the corresponding Lagrange multiplier.
Proof.
For energies, we choose a text function $\rho_{R}(x)=R^{-3}\rho(R^{-1}x)\in W_{R}$ for some suitable function $\rho$, and estimate $E\left(\rho_{R}\right)=o\left(R^{-1}\right)-cR^{-1}$ as $R\rightarrow\infty$, see [4, Lemma 6], then $\exists R_{2}$ such that $\forall R\geq R_{2}$, $E\left(\sigma_{R}\right)\leq E\left(\sigma_{R_{2}}\right)\leq E\left(\rho_{R_{2}}\right)\coloneq\alpha<0$.
For Lagrange multipliers, by Hölder’s inequality, we have
$\alpha\leq-E\left(\sigma_{R}\right)\leq\frac{1}{2}\int_{\mathbb{R}^{3}}\sigma_{R}V_{\sigma_{R}}\,dx\leq\frac{1}{2}m\left\|V_{\sigma_{R}}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}$
then for $R$ large enough, $\left\|V_{\sigma_{R}}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\geq-\frac{2\alpha}{m}$, which can help to imply $\sigma_{R}$ cannot become uniformly diffuse as $R\rightarrow\infty$. That is, we claim $\exists\epsilon_{0}$, such that $\forall R>0,\exists x_{R}\in\mathbb{R}^{3}$, we have
$\int_{\left|y-x_{R}\right|<1}\sigma_{R}(y)\,dy\geq\epsilon_{0}$
In fact, let $\epsilon_{R}=\sup\limits_{x}\int_{|y-x|<1}\rho_{R}\,dy$, and estimate $V_{\sigma_{R}}(x)$ as
$V_{\sigma_{R}}(x)$ $=\int_{|y-x|<1}\frac{\sigma_{R}(y)}{|y-x|}\,dy+\int_{1<|y-x|<r}\frac{\sigma_{R}(y)}{|y-x|}\,dy+\int_{|y-x|>r}\frac{\sigma_{R}(y)}{|y-x|}\,dy$
$=V_{1}+V_{2}+V_{3}$
It’s easy to see $V_{3}\leq\frac{M}{r}$. Since the shell 1<|yx|<r1<|y-x|<r can be covered by finite balls of radius 1 (it is 3-dimensional space thus the number of balls can be $\leq cr^{3}$), then $V_{2}\leq c\epsilon_{R}r^{3}$. Thanks to Lemma 2.21 $\left(\left\|\sigma_{R}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\leq k_{1}\right)$ and Proposition A.3, we have $V_{1}\leq c\left(\epsilon_{R}^{a}+\epsilon_{R}^{b}\right)$ where $a>0,b>0$. Let $g(r)=\frac{M}{r}+c\epsilon_{R}r^{3}+c\left(\epsilon_{R}^{a}+\epsilon_{R}^{b}\right)$, where $r>0$. We know it has minimal value $g_{\min}=\widetilde{c}\epsilon_{R}^{\frac{1}{4}}+c\left(\epsilon_{R}^{a}+\epsilon_{R}^{b}\right)$ at $r_{\min}=\left(\frac{M}{3c\epsilon_{R}}\right)^{\frac{1}{4}}$. If $r_{\min}\geq 1$, then $\left\|V_{\sigma_{R}}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\leq\widetilde{c}\epsilon_{R}^{\frac{1}{4}}+c\left(\epsilon_{R}^{a}+\epsilon_{R}^{b}\right)$. On the other hand, $\left\|V_{\sigma_{R}}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\geq-\frac{2e}{m}$. Therefore, $\exists\widetilde{\epsilon}>0,\widetilde{R}>0$, such that $\forall R>\widetilde{R}$, we have $\epsilon_{R}>\widetilde{\epsilon}$. If $r_{\text{min }}<1$, then $\epsilon_{R}>\frac{M}{3c}$, with an abuse of notation, $\epsilon_{R}>\widetilde{\epsilon}$ still holds by letting $\widetilde{\epsilon}=\min\left\{\widetilde{\epsilon},\frac{M}{3c}\right\}$. If $R\leq\widetilde{R},\sigma_{R}\in W_{\widetilde{R}}$, thus spt $\sigma_{R}$ is covered by finite balls with radius 1 (the number of balls $C$ is independent of $R$). Since the mass of $\sigma_{R}$ is $m$, we know $\epsilon_{R}>\frac{m}{2C}$. Let $\epsilon=\min\left\{\widetilde{\epsilon},\frac{m}{2C}\right\}$, we know $\epsilon_{R}>\epsilon$ for all $R$, which proves the claim.
By the construction of $W_{R}$ we get $\sigma_{R}$ is axially symmetric, then $x_{R}$ mentioned in the claim should satisfy $r\left(x_{R}\right)\leq r_{0}$.
Therefore, we can modify the arguments to prove $\lambda_{R}<0$ in Step (1) by choosing suitable $x_{0}\in B_{r}\left(x_{R}\right)$ to get $\lambda_{R}\leq l<0$ for all large $R$. ∎
Remark 2.25.
One can use an argument similar to the proof of Lemma 2.24 above to show that, in fact, for any energy minimizer, the corresponding conclusion holds. That is, the energy and the Lagrangian multiplier are both negative.
Remark 2.26.
The result that the energy is negative is nice, since later we will show $\sigma_{R}$ is truly a minimizer on $W_{\infty}=mR{(\mathbb{R}^{3})}$, then $E_{0}\left(\sigma_{R}\right)$ is the minimal energy $e_{0}(m)$, which is less than 0.
Lemma 2.27 (Uniform Bound on the Sizes of Supports [4]).
$\exists R_{1}>0$, such that $\forall R>R_{1}$, spt $\sigma_{R}$ is contained in $\overline{B_{R_{1}}(0)}$.
Proof.
We first show $V_{\sigma_{R}}(x)\leq-l$ holds when the distance from $x$ to the z-axis $r(x)$ and $R$ is large enough, then due to (23) and Lemma 2.24 we know $\sigma_{R}(x)=0$. Let $a,b>0$ be as in the estimate $V_{1}\leq c(\epsilon_{R}^{a}+\epsilon_{R}^{b})$ from the proof of Lemma 2.24, and let $c_{1},c_{2},c_{3}>0$ be sufficiently large constants. Then define
$f(r,\pi)=c_{1}(\pi^{-a}+\pi^{-b})+c_{2}\pi^{-1}r^{3}+c_{3}r^{-1}$
Notice if we choose $r\geq r_{0}:=1+\frac{-3c_{3}}{l}$, then $c_{3}r^{-1}<-\frac{l}{3}$, and we have $f(r,\pi)\leq-l$ when $\pi$ is large enough.
Now we claim $\exists\pi_{0}>2r_{0}-1$, $\forall\pi>\pi_{0}$, $\forall x$ with $r(x)>\pi$, $\forall r\in\left(r_{0},\frac{1+\pi}{2}\right)$ we have $V_{\sigma_{R}}(x)\leq f(r,\pi)\leq-l$.
In fact, let $\epsilon_{\pi,R}=\sup\limits_{r(x)>\frac{\pi}{2}}\int_{|y-x|<1}\sigma_{R}(y)\,dy$, then $\exists x_{\pi,R}\in\mathbb{R}^{3}$ with $r\left(x_{\pi,R}\right)>\frac{\pi}{2}$, such that
$\int_{\left|y-x_{\pi,R}\right|<1}\sigma_{R}(y)\,dy>\epsilon_{\pi,R}-\frac{1}{\pi}$
On the other hand, due to the axial symmetry of $\sigma_{R}(y),\sigma_{R}(y)$ has mass $\geq{cr}\left(x_{\pi,R}\right)\left(\epsilon_{\pi,R}-\frac{1}{\pi}\right)$ in the torus obtained by revolving the ball $\left|y-x_{\pi,R}\right|<1$ around the z-axis, which implies $\epsilon_{\pi,R}\leq\frac{m}{{cr}\left(x_{\pi,R}\right)}+\frac{1}{\pi}<\frac{2m}{c\pi}+\frac{1}{\pi}$. By choosing $c$ sufficiently large, we have $\epsilon_{\pi,R}<\frac{c}{\pi}$.
For all $x$ with $r(x)>\pi$, notice due to $r<\frac{1+\pi}{2}$, we can choose finite balls centered at $x_{n}$ of radius 1 with $r\left(x_{n}\right)>\frac{\pi}{2}$ covering 1<|yx|<r1<|y-x|<r, and the number of balls can be $\leq cr^{3}$. Similar to Lemma 2.24, together with $\epsilon_{\pi,R}<\frac{c}{\pi}$, we have
$V_{\sigma_{R}}(x)$ $=\int_{|y-x|<1}\frac{\sigma_{R}(y)}{|y-x|}\,dy+\int_{1<|y-x|<r}\frac{\sigma_{R}(y)}{|y-x|}\,dy+\int_{|y-x|>r}\frac{\sigma_{R}(y)}{|y-x|}\,dy$
$\leq c\left(\epsilon_{\pi,R}^{a}+\epsilon_{\pi,R}^{b}\right)+cr^{3}\epsilon_{\pi,R}+cr^{-1}$
$\leq c_{1}\left(\pi^{-a}+\pi^{-b}\right)+c_{2}\pi^{-1}r^{3}+c_{3}r^{-1}$
$=f(r,\pi)$
Since $r\geq r_{0}$, when $\pi$ is large enough we have $f(r,\pi)\leq-l$ as we mention above. Therefore, the claim holds true. Thus $\forall\pi>\pi_{0}$, $\forall x$ with $r(x)>\pi$, we have $V_{\sigma_{R}}(x)\leq f(\frac{2r_{0}+1+\pi}{4},\pi)\leq-l$, then $\sigma_{R}(x)=0$ due to (23) and Lemma 2.24.
Now we consider the z-direction and show $V_{\sigma_{R}}(x)\leq-l$ holds when $z(x)$ is large enough, which is relatively easier. We can first show $\exists\delta>0$, $r\geq 2$ and $\widetilde{R_{0}}>0$, such that $\forall R\geq\widetilde{R_{0}}$, if $\int_{|z(x)-d|<r}\sigma_{R}(x)\,dx<\delta$, then $\sigma_{R}(x)=0$ for $|z(x)-d|\leq 1$. This comes from that when $|z(x)-d|\leq 1$ we have the estimate:
$V_{\sigma_{R}}(x)=\int_{|z(y)-d|<r}\frac{\sigma_{R}(x)}{|y-x|}\,dy+\int_{|z(y)-d|>r}\frac{\sigma_{R}(x)}{|y-x|}\,dy\leq c\left(\delta^{a}+\delta^{b}\right)+\frac{m}{r-1}$
And then letting $\delta$ small enough and $r$ large enough can give $V_{\sigma_{R}}(x)\leq-l$, which means $\sigma_{R}(x)=0$. Then we set $Z_{n}=\{x||z(x)-2n\mid\leq 1\},n=0,\pm 1,\pm 2,\ldots$. and $Z_{n}^{\prime}=\{x||z(x)-2n\mid<r\}$, without loss of generality, we can take $r$ to be an integer. Let $N$ be the number of $Z_{n}^{\prime}$ such that $\int_{Z_{n}^{\prime}}\sigma_{R}\,dx\geq\delta$. Since the mass of $\sigma_{R}$ is $m<\infty$ and the $Z_{n}^{\prime}$s overlap finite times $\widetilde{r}$$\infty$, we have $N\delta\leq\widetilde{r}m$, i.e., $N\leq\frac{\widetilde{r}m}{\delta}$. Therefore, $\sigma_{R}(x)=0$ for large $R$ except on a set of horizontal slabs $Z_{n}$ of total height at most $\frac{2\widetilde{r}m}{\delta}$.
If these $Z_{n}{s}$ were known to be contiguous, then we would have $\sigma_{R}(x)=0$ for $|z(x)|\geq\frac{2rm}{\delta}$. But if the $Z_{n}{s}$ were not contiguous, we could follow the idea of the strong rearrangement inequality in Lieb [43, Lemma 3]. We could rearrange the order of $Z_{n}$, slide the half-space above an empty $Z_{n}$ down one unit and obtain a new $\widetilde{\sigma_{R}}\in W_{R}$ from $\sigma_{R}$. During the process we notice
$U\left(\widetilde{\sigma_{R}}\right)=\int_{\mathbb{R}^{3}}A\left(\widetilde{\sigma_{R}}\right)\,dx=\int_{\mathbb{R}^{3}}A\left(\sigma_{R}\right)\,dx=U\left(\sigma_{R}\right)$
$G\left(\widetilde{\sigma_{R}},\widetilde{\sigma_{R}}\right)=\int_{\mathbb{R}^{3}}\widetilde{\sigma_{R}}V_{\widetilde{\sigma_{R}}}\,dx>\int_{\mathbb{R}^{3}}\sigma_{R}V_{\sigma_{R}}\,dx=G\left(\sigma_{R},\sigma_{R}\right)$
then we obtain $E_{0}\left(\widetilde{\sigma_{R}}\right)<E_{0}\left(\sigma_{R}\right)$, which leads to a contradiction.
Collect the results above, we know we can choose a large $R_{1}$, such that $\forall R>R_{1}$, spt $\sigma_{R}$ is contained in $\overline{B_{R_{1}}(0)}$. ∎
Now we can go forward to Step (4) and prove the constrained minimizer can be a global minimizer.
Lemma 2.28 (Existence of Global Minimizer in Non-rotating case [4]).
Take $R^{\prime}=\max\left\{k_{1},R_{1}\right\}+1$, where $k_{1}$ and $R_{1}$ are given in Lemma 2.21 and Lemma 2.27, then we have $\forall R\geq R^{\prime},\sigma_{R}$ is a minimizer on $m{R}\left(\mathbb{R}^{3}\right)$.
Proof.
Thanks to Lemma 2.21 and Lemma 2.27, we know $\forall R\geq R^{\prime},\sigma_{R}\in W_{R^{\prime}}$, which implies $E_{0}\left(\sigma_{R}\right)=E_{0}\left(\sigma_{R^{\prime}}\right)$ since they should both be minimizers on $W_{R^{\prime}}$. Thus, we just need to show $\sigma_{R^{\prime}}$ is a minimizer on $W_{\infty}=m{R}\left(\mathbb{R}^{3}\right)$. To show that, given $\rho\in W_{\infty}$, we construct a $\rho_{R}\in W_{R}$ for large $R$, with $E_{0}\left(\rho_{R}\right)\leq E_{0}(\rho)+\epsilon$, where $\epsilon$ can be arbitrarily small. (For the construction of $\rho_{R}$, although $\rho$ is not necessarily to be spherically symmetric, but one can apply a strong rearrangement inequality in Lieb [43, Lemma 3] and the generalization by Sobolev of the Hardy-Littlewood theorem on rearrangements of functions [54] to replace $\rho$ by another spherically symmetric and radially decreasing (after translation) $\rho$, see for example [4, Appendix], [46, Section 6]. Then one can refer to [4, Section 6] or [16, Section 4] to construct $\rho_{R}$.) Then we have $E_{0}\left(\sigma_{R^{\prime}}\right)=E_{0}\left(\sigma_{R}\right)\leq E_{0}\left(\rho_{R}\right)\leq E_{0}(\rho)+\epsilon$. Let $\epsilon\rightarrow 0$ we obtain $E_{0}\left(\sigma_{R^{\prime}}\right)\leq E_{0}(\rho)$, which shows the result. ∎
It is natural to conjecture whether there are some relations between Lagrange-multipliers $\lambda_{m}$ and the derivatives of $e_{0}(m)$, since they can be obtained by somehow “differentiating” $E_{0}\left(\sigma_{m}\right)$. Thanks to Lieb and Yau’s paper [42, Theorem 3 and Theorem 4], we can know the left and right derivatives of $e_{0}(m)$ bound $\lambda:e_{0}^{\prime}\left(m^{+}\right)\leq\lambda\leq e_{0}^{\prime}\left(m^{-}\right)$, as cited in [46, Theorem 3.5] and [17, Section 2]. Moreover, we find a more precise quantitative property in the case that complies with polytropic equations of state in the following.
Proposition 2.29 (Expression for Lagrange-multipliers).
Let pressure $P(\rho)$ satisfies polytropic law with index $\gamma>\frac{4}{3}$, $E_{0}(\rho)$ from (5), $e_{0}(m)$ from (12) and $m\in[0,1]$, then $e_{0}^{\prime}(m)=\lambda_{m}=-(5\gamma-6)m^{\frac{2\gamma-2}{3\gamma-4}}U(\sigma)<0$, where $\sigma$ is the minimizer of $E_{0}(\rho)$ with mass $1$.
Proof.
By scaling relations in Section 3 (Theorem 3.2) we can obtain $e_{0}(m)=m^{\frac{5\gamma-6}{3\gamma-4}}e_{0}$, which implies $e_{0}^{\prime}$ exists. Then $e_{0}^{\prime}(m)=\lambda_{m}$ can come from Theorem 2.6 (viii). As in the proof of Theorem 2.6 (ii), we use the scaling relations mentioned in Section 3 (Proposition 3.4). We obtain $E_{0}^{\prime}\left(\sigma_{m}\right)(x)=\frac{1}{A^{\gamma-1}}E_{0}^{\prime}(\sigma)\left(\frac{1}{B}x\right)$, thus $\lambda_{m}=\frac{1}{A^{\gamma-1}}\lambda_{1}$. Since $e_{0}(m)=m^{\frac{5\gamma-6}{3\gamma-4}}e_{0}$, then $e_{0}^{\prime}(m)=\frac{5\gamma-6}{3\gamma-4}m^{\frac{2\gamma-2}{3\gamma-4}}e_{0}$, and $e_{0}^{\prime}(m)=\frac{1}{A^{\gamma-1}}e_{0}^{\prime}(1)$. Therefore, we just need to consider the case $m=1$. Notice we have Euler-Lagrange equation $E_{0}^{\prime}(\sigma)(x)=\lambda_{1}$ a.e. on $\{\sigma(x)>0\}$. Rather than substituting a specific $x$ to calculate $E_{0}^{\prime}(\sigma)$, we view $E_{0}^{\prime}(\sigma)$ as a functional in the sense that $E_{0}^{\prime}(\sigma)(\rho)=\int_{\mathbb{R}^{3}}E_{0}^{\prime}(\sigma)\rho\,dx$ as in Lemma 2.9 or Remark 2.10. We take $\rho=\sigma$ as a test function, notice the mass of $\sigma$ is 1. Then we have
$\int_{\mathbb{R}^{3}}E_{0}^{\prime}(\sigma)\sigma\,dx=\int_{\{\sigma(x)>0\}}E_{0}^{\prime}(\sigma)\sigma\,dx=\lambda_{1}\int_{\{\sigma(x)>0\}}\sigma\,dx=\lambda_{1}$
On the other hand, we have
$\int_{\mathbb{R}^{3}}E_{0}^{\prime}(\sigma)\sigma\,dx=\int_{\mathbb{R}^{3}}\left(A^{\prime}(\sigma)-V_{\sigma}\right)\sigma\,dx=\gamma U(\sigma)-G(\sigma,\sigma)$
Then we get $\lambda_{1}=\gamma U(\sigma)-G(\sigma,\sigma)$, while $e_{0}^{\prime}(1)=\frac{5\gamma-6}{3\gamma-4}e_{0}=\frac{5\gamma-6}{3\gamma-4}\left(U(\sigma)-\frac{G(\sigma,\sigma)}{2}\right)$. Due to the previous arguments that $\lambda_{1}=e_{0}^{\prime}(1)$, we can deduce that $\lambda_{1}=e_{0}^{\prime}(1)=(6-5\gamma)U(\sigma)$ if $\gamma\neq 2$. In the case $\gamma=2$, we can apply Remark 2.39 below and obtain $\frac{G(\sigma,\sigma)}{2}=(3\gamma-3)U(\sigma)$, then we also have $\lambda_{1}=e_{0}^{\prime}(1)=(6-5\gamma)U(\sigma)$. ∎
2.3 Uniqueness Results for Non-rotating Bodies
In this subsection, we discuss uniqueness results. To show the uniqueness result in the following, we follow Lieb and Yau’s arguments [42, Lemma 10, Lemma 11 and remark following] [46], but adapt it from the framework of quantum mechanics to that of classical mechanics. We show if, in addition, $P(\rho)$ satisfies (F4), $A^{\prime}\left(\rho^{3}\right)$ has second derivative and is convex, uniqueness of minimizer up to translation can also be shown in the following arguments. In particular, if polytropic law holds true with $\gamma>\frac{4}{3}$, then $A^{\prime}\left(\rho^{3}\right)=\frac{K\gamma}{\gamma-1}\rho^{3(\gamma-1)}$ satisfies those conditions.
Due to Theorem 2.6 (iii), we know the minimizer $\sigma_{m}$ is spherically symmetric after translation. In general, given a spherically symmetric function $\rho$, with an abuse of notation, we denote $\rho(|x|)=\rho(s)$, and
$m(r)=m_{\rho}(r)=\int_{|x|<r}\rho\,dx=4\pi\int_{0}^{r}\rho(s)s^{2}ds$
is the mass of $\rho$ in $\{|x|<r\}$, then easy to show (we assume the following computational results are finite. This condition holds, for instance, when $\rho$ is continuous with compact support):
$V_{\rho}(x)=V_{\rho}(|x|)=\frac{m(|x|)}{|x|}+4\pi\int_{|x|}^{\infty}t\rho(t)dt$ (26)
By direct computation we have
$\frac{d}{dr}V_{\rho}(r)=\frac{-m(r)}{r^{2}},\text{ where }r\neq 0$ (27)
$\nabla V_{\rho}(x)=\frac{d}{dr}V_{\rho}(|x|)\cdot\frac{x}{|x|}=\frac{-m(|x|)}{|x|^{2}}\cdot\frac{x}{|x|},\text{ where }x\neq 0$ (28)
Note that $V_{\rho}(0)=V_{\rho}(|0|)=4\pi\int_{0}^{\infty}t\rho(t)dt$, which follows directly from the definition of $V_{\rho}$ (or by applying L’Hôpital’s rule to show $V_{\rho}(0)=\lim\limits_{x\rightarrow 0}V_{\rho}(x)=4\pi\int_{0}^{\infty}t\rho(t)dt$). Then after using L’Hôpital’s rule [52, Theorem 5.13], one can also show that
$\frac{d}{dr}V_{\rho}(0)=0$ (29)
$\nabla V_{\rho}(0)=0$ (30)
Another way to show $\nabla V_{\rho}(0)=0$ is the fact that since $\lim\limits_{x\rightarrow 0}\nabla V_{\rho}(x)=0$ exists, by mean value theorem we can prove $\nabla V_{\rho}(0)=\lim\limits_{x\rightarrow 0}\nabla V_{\rho}(x)=0$. The third way to show it is using the fact $V_{\rho}$ is also spherically symmetric (same for $\frac{d}{dr}V_{\rho}(0)$) [^5] Strictly speaking, we need to verify that $V$ is differentiable at 0 then use the symmetric property. This differentiability can be deduced from the regularity of $\rho$. For instance, if $\rho$ is continuous with compact support, then arguments similar to those in Theorem 2.6 (vi) ensure the differentiability of $V$ at 0..
Based on (26), we can show the Shell Theorem, which is given by Newton.
Theorem 2.30 (Shell Theorem [57] [49, Page 192 Theorem XXX and Page 193 Theorem XXXI] [48, Page 218 Theorem XXX and Theorem XXXI]).
Suppose $\rho$ is spherically symmetric, if $\rho$ is supported in $B_{R}(0)$, denote its total mass by $m=\int_{0}^{R}4\pi r^{2}\rho(r)\,dr$, then the gravitational potential $V_{\rho}$ satisfies:
  1. (i)
    For exterior points ($|x|\geq R$): $V_{\rho}(x)=\frac{m}{|x|}$. That is, given an object $P$ outside $B_{R}(0)$, $\rho$ produces the same external gravitational potential as if all $\rho$’s mass were concentrated at its center.
  2. (ii)
    For interior points ($|x|\leq R$): $V_{\rho}(x)=\frac{m(|x|)}{|x|}+4\pi\int_{|x|}^{\infty}t\rho(t)dt$, where the second term shows that the outer shells contribute a constant potential inside. In particular, if $\rho$ vanishes in $B_{r}(0)$ for some $r<R$, then $V_{\rho}$ is a constant in $B_{r}(0)$. Given an object $P$ inside $B_{r}(0)$, no net gravitational force is exerted by $\rho$ on $P$ due to (27).
Remark 2.31.
In fact, one can also utilize Legendre polynomials to prove the Shell Theorem Theorem 2.30. The idea is essentially the same as the proof of Corollary 2.32 in the following.
Corollary 2.32 (Gravitational Potential Energy between Two Spherical Bodies).
Suppose $\rho$ and $\sigma$ are spherically symmetric with total masses $m_{1}$ and $m_{2}$, and their centers are separated by a distance $R$. If $R$ is greater than the sum of their radii (so that the spheres do not overlap), then their mutual gravitational potential energy $G(\rho,\sigma)$ is
$G(\rho,\sigma)=\frac{m_{1}m_{2}}{R}$
In other words, two non-overlapping spherically symmetric bodies attract each other exactly as if all their mass were concentrated at their centers.
Proof.
Without loss of generality, we can set up a coordinate system such that the center of $\rho$ is $(0,0,0)$, and the center of $\sigma$ is $\mathbf{D}=(0,0,R)$. Assume the radius of $\rho$ is $r_{1}$ and the radius of $\sigma$ is $r_{2}$, notice $R\geq r_{1}+r_{2}$, then thanks to Theorem 2.30 the mutual gravitational potential energy is given by:
$G(\rho,\sigma)=\int_{B_{r_{1}}(0)}\rho(x)\,V_{\sigma}(x)\,dx=m_{2}\int_{B_{r_{1}}(0)}\frac{\rho(x)}{|x-\mathbf{D}|}\,dx$ (31)
We can expand $1/|x-\mathbf{D}|$ as:
$\frac{1}{|x-\mathbf{D}|}=\frac{1}{R}\sum_{\ell=0}^{\infty}\left(\frac{|x|}{R}\right)^{\ell}P_{\ell}(\cos\gamma)$ (32)
where $P_{\ell}$ is the Legendre’s polynomial, $\gamma$ is the angle between $x$ and $\mathbf{D}$.
Then plug (32) into (31), and notice $\rho$ is spherically symmetric, which implies the Legendre polynomials expansion of $\rho$ is itself — $\rho(x)=\rho(|x|)$), then we know all terms involving $P_{\ell}(\cos\gamma)$ for $\ell>0$ vanish upon angular integration, leaving only the $\ell=0$ term:
$G(\rho,\sigma)=m_{2}\int_{B_{r_{1}}(0)}\frac{\rho(x)}{|x-\mathbf{D}|}\,dx=\frac{m_{2}}{R}\int_{B_{r_{1}}(0)}\rho(x)\,dx=\frac{m_{1}m_{2}}{R}$
This is the result we want. ∎
Remark 2.33.
Notice that in Corollary 2.32, it is not necessary to assume constant density, i.e., the celestial body need not be homogeneous. In the case of a homogeneous celestial body, we may alternatively use the fact that $\frac{1}{|x-\mathbf{D}|}$ is a harmonic function in $B_{r_{1}}(0)$ and apply the mean value property of harmonic function to prove Corollary 2.32.
Let’s recall that for minimizer $\sigma$, by Theorem 2.6 (vii) we have the Euler-Lagrange equation (EL):
$A^{\prime}(\sigma(x))=\left[V_{\sigma}(x)+\lambda\right]_{+}$ (EL)
In particular,
$E_{0}^{\prime}(\sigma)(x)=\lambda\text{, where }\sigma(x)>0$ (33)
More specifically,
$E_{0}^{\prime}(\sigma)(r)=A^{\prime}(\sigma(r))-V_{\rho}(r)=\lambda\text{, where }\sigma(r)>0$ (34)
Due to (26), after taking the derivative of (34), we can have
$\frac{d}{dr}A^{\prime}(\sigma(r))=-r^{-2}m(r)\text{, where }\sigma>0$ (35)
Hence $\sigma$ is decreasing when it is positive. Moreover, from (33) we also have $\sigma(x)=\left(A^{\prime}\right)^{-1}\left(V_{\sigma}(x)+\lambda\right)$, notice that $\Delta\left(V_{\sigma}+\lambda\right)=\Delta V_{\sigma}=-4\pi\sigma$, set $\Theta_{\sigma}=V_{\sigma}+\lambda$, then we have
$\Delta\Theta_{\sigma}=4\pi\sigma=4\pi\left(A^{\prime}\right)^{-1}\left(\Theta_{\sigma}\right)\text{, where }\sigma>0$ (36)
We relax the constraints and allow the equation to hold over the entire $\mathbb{R}^{3}$.
$\Delta\Theta_{\sigma}=4\pi\left(A^{\prime}\right)^{-1}\left(\Theta_{\sigma}\right)$ (37)
Remark 2.34.
As mentioned in [42], since $\sigma$ is spherically symmetric, so is $\Theta_{\sigma}$. Thus (36) and (37) turn out to be a second order ODE. Due to (30) we have $\nabla\Theta_{\sigma}(0)=0$. If $\Theta_{\sigma}(0)=\beta>0$, by uniqueness theorem of ODE (for example, Picard’s method), we know (37) has a unique positive solution up to some $R(\beta)$ and spt $\Theta_{\sigma}=\overline{B_{R(\beta)}(0)}$. Since $\Delta\Theta_{\sigma}=-4\pi\left(A^{\prime}\right)^{-1}\left(\Theta_{\sigma}\right)=-4\pi\sigma$, we can see $B_{R(\beta)}(0)=\{\sigma>0\}$. Therefore, $0<R(\beta)<\infty$ due to Theorem 2.6 (v).
On the other hand, thanks to Lieb and Yau [42, Lemma 8], we can show $0<R(\beta)<\infty$ holds true for any radial solutions to (37) under certain assumptions of initial conditions mentioned above and of pressure [42]. Furthermore, note that our preceding arguments are not restricted to a specific mass (e.g., $m=1$), but naturally generalizes to any mass $m>0$. Consequently, the positive solutions to (37) actually correspond one-to-one with the solutions to (EL). We will return to discuss more about these relations later.
Although the solutions to (EL) can correspond one-to-one with the positive solutions to (37), and one positive solution to (37) is unique given initial conditions by uniqueness theorem of ODE, so far we cannot conclude that the minimizer is unique. The problem is we do not yet know two minimizers with same mass have to correspond with the same initial conditions of (37). The next several results help to show the minimizer is truly unique.
Lemma 2.35 (Relation between Central Density and Mass [42, Lemma 10]).
Suppose $\sigma_{m}$ and $\sigma_{\widetilde{m}}$ are minimizers for $E_{0}(\rho)$ with $\int_{\mathbb{R}^{3}}\sigma_{m}\,dx=m$, $\int_{\mathbb{R}^{3}}\sigma_{\widetilde{m}}\,dx=\widetilde{m}$ respectively and their centers of mass are $(0,0,0)^{T}$. Let $R_{m}$ and $R_{\widetilde{m}}$ be the radii of their supports and set $R=\max\left\{R_{m},R_{\widetilde{m}}\right\}$. Suppose that $\sigma_{m}(0)>\sigma_{\widetilde{m}}(0)$, then $\forall 0<r<R$ we have $m(r)>\widetilde{m}(r)$.
Proof.
We provide a proof outline here. One can check the proofs in [42] and references therein, to understand some step(s) omitted in the following.
[…]
We consider dividing the density functions into inner parts and outer parts. That is, let
$\sigma_{m}^{i}=\sigma_{m}\mathbf{1}_{\{|x|\leq\bar{r}\}},\quad\sigma_{m}^{o}=\sigma_{m}\mathbf{1}_{\{|x|>\bar{r}\}},$
$\sigma_{\widetilde{m}}^{i}=\sigma_{\widetilde{m}}\mathbf{1}_{\{|x|\leq\bar{r}\}},\quad\sigma_{\widetilde{m}}^{o}=\sigma_{\widetilde{m}}\mathbf{1}_{\{|x|>\bar{r}\}}.$
Set $E_{0}^{i}\left(\rho^{i}\right)=E_{0}\left(\rho^{i}\right)$, and
$E_{0}^{o}\left(\rho^{o}\right)=U\left(\rho^{o}\right)-\frac{G\left(\rho^{o},\rho^{o}\right)}{2}-Q\int_{\mathbb{R}^{3}}\frac{\rho^{o}(x)}{|x|}\,dx$
By Remark 2.30, $\int_{\mathbb{R}^{3}}\sigma_{m}^{i}\,dx=\int_{\{|x|\leq\widetilde{r}\}}\sigma_{m}^{i}\,dx=Q$, and notice
$G\left(\sigma_{m}^{i},\sigma_{m}^{o}\right)$ $=\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\frac{\sigma_{m}^{i}(y)\sigma_{m}^{o}(x)}{|x-y|}\,dxdy$
$=\int_{\{|y|\leq\widetilde{r}\}}\sigma_{m}^{i}(y)V_{\sigma_{m}^{o}}(y)\,dy$
[…]
$=QV_{\sigma_{m}^{o}}(0)$
$=Q\int_{\mathbb{R}^{3}}\frac{\sigma_{m}^{o}(x)}{|x|}\,dx.$
we obtain $e_{0}(m)=E_{0}\left(\sigma_{m}\right)=E_{0}^{i}\left(\sigma_{m}^{i}\right)+E_{0}^{o}\left(\sigma_{m}^{o}\right)$. Similarly we have $e_{0}(\widetilde{m})=E_{0}\left(\sigma_{\widetilde{m}}\right)=E_{0}^{i}\left(\sigma_{m}^{i}\right)+E_{0}^{o}\left(\sigma_{\widetilde{m}}^{o}\right)$.
Moreover, we claim $E_{0}^{i}\left(\sigma_{m}^{i}\right)=E_{0}^{i}\left(\sigma_{\widetilde{m}}^{i}\right)={\inf\limits_{\rho^{i}\in U^{i}}{E_{0}^{i}\left(\rho^{i}\right)}}$, where
$U^{i}=\left\{\rho^{i}\in L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)\middle|\rho^{i}\geq 0,\rho^{i}\text{ vanishes outside }\overline{B_{\widetilde{r}}(0)},~{\int_{\mathbb{R}^{3}}\rho^{i}\,dx}=Q\right\}$
[…]
Remark 2.36.
Notice that in Lemma 2.35, only when $\forall 0<r<R$ we have $m(r)>\widetilde{m}(r)$. Hence it does not say $m>\widetilde{m}$, although we shall later see that this is true (Corollary 2.43 or Remark 3.3) after obtaining the uniqueness result. If we knew in advance that $m>\widetilde{m}$, the proof of the uniqueness of minimizer would be trivial.
So far we only know the energy consists of inertial energy and gravitational interaction energy:
$E_{0}(\rho)=U(\rho)-\frac{G(\rho,\rho)}{2}$ (38)
To find a deeper relation between energy and $A(\rho)$, it will be nice if we can find another expression of $E_{0}(\rho)$ in which $G(\rho,\rho)$ does not appear. Thanks to Lieb and Yau’s arguments [42], we can truly find such expression when $\rho$ is a minimizer, since in this case we have Euler-Lagrange equation which connects $A^{\prime}(\rho)$ with gravitational potential $V_{\rho}$ and thereby $G(\rho,\rho)$.
Lemma 2.37 (Relation between $E_{0}$ and $A$ [42, Lemma 6]).
Let $P(s)$ satisfies (F1) (F2) (F3). Suppose $\sigma$ is continuous, spherically symmetric with finite mass $m=\int_{\mathbb{R}^{3}}\sigma\,dx$, $\sigma$ has compact support, and satisfies (EL) on all of $\mathbb{R}^{3}$ for $J=0$ and a single $\lambda<0$. Then
$E_{0}(\sigma)=\int_{\mathbb{R}^{3}}\left(4A(\sigma)-3\sigma A^{\prime}(\sigma)\right)\,dx$ (39)
Proof.
If $\sigma=0$, then (39) is trivial since both sides are 0. If $\sigma$ is not zero function, since $\sigma\in L^{1}\left(\mathbb{R}^{3}\right)\cap C_{c}^{0}\left(\mathbb{R}^{3}\right)$, by Proposition A.4 we know $V_{\rho}$ is continuously differentiable, thus we can take the derivative of (34) where $\sigma>0$ (notice $\{\sigma>0\}$ is open). Since $\sigma$ is spherically symmetric, we have exactly (35):
$\frac{d}{dr}A^{\prime}(\sigma(r))=-r^{-2}m(r)$
Multiply it by $4\pi r^{3}\sigma(r)$ and integrate over $\mathbb{R}$, we obtain
$4\pi\int_{0}^{\infty}r^{3}\left(\frac{d}{dr}A^{\prime}(\sigma)\right)\sigma dr=4\pi\int_{0}^{\infty}m(r)\sigma(r)rdr$
Easy to see $\iint_{\{|x|\leq|y|\}}\frac{\sigma(x)\sigma(y)}{|x-y|}\,dx\,dy=\iint_{\{|x|\geq|y|\}}\frac{\sigma(x)\sigma(y)}{|x-y|}\,dx\,dy$. Then we compute
$\frac{G(\sigma,\sigma)}{2}$ $=\frac{1}{2}\left(\iint_{\{|x|\leq|y|\}}\frac{\sigma(x)\sigma(y)}{|x-y|}\,dx\,dy+\iint_{\{|x|\geq|y|\}}\frac{\sigma(x)\sigma(y)}{|x-y|}\,dx\,dy\right)$
$=\iint_{\{|x|\leq|y|\}}\frac{\sigma(x)\sigma(y)}{|x-y|}\,dx\,dy$
$=\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\frac{\sigma(x)\sigma(y)\mathbf{1}_{\{|x|\leq|y|\}}}{|x-y|}\,dx\,dy$
$=\int_{\mathbb{R}^{3}}\sigma(y)\,dy\int_{\mathbb{R}^{3}}\frac{\sigma(x)\mathbf{1}_{\{|x|\leq|y|\}}}{|x-y|}\,dx$
$=\int_{\mathbb{R}^{3}}\sigma(y)V_{\left\{\sigma\mathbf{1}_{\{|\cdot|\leq|y|\}}\right\}}(y)\,dy$
[…]
$=4\pi\int_{0}^{\infty}m(r)\sigma(r)rdr$
Notice we use spherical coordinate transformation in the last identity.
On the other hand, we notice
$\left(\frac{d}{dr}A^{\prime}(\sigma)\right)\sigma=\frac{d}{dr}\left(\sigma A^{\prime}(\sigma)-A(\sigma)\right)$
And $\sigma$ vanishes at 0 and outside a large ball, so is $\sigma A^{\prime}(\sigma)-A(\sigma)$. Then we use spherical coordinate transformation and integration by parts, and obtain
[…]
$=-4\pi\int_{0}^{\infty}r^{3}\left(\frac{d}{dr}A^{\prime}(\sigma)\right)\sigma dr$
Collect the results above we obtain $E_{0}(\sigma)=\int_{\mathbb{R}^{3}}A(\sigma)\,dx-\frac{G(\sigma,\sigma)}{2}=\int_{\mathbb{R}^{3}}(4A\left(\sigma)-3\sigma A^{\prime}(\sigma)\right)\,dx$. ∎
Remark 2.38.
Although the differentiability of $\sigma$ is used in Lieb and Yau’s proof [42], we refine their proof as above and do not necessarily need the differentiability of $\sigma$. This is the reason we do not need to assume (F4) holds. If $P(s)$ satisfies (F1)(F2)(F3), then thanks to Theorem 2.6, a minimizer $\sigma_{m}$ (after translation) meets the conditions in Lemma 2.37, thus (39) holds true for $\sigma_{m}$.
Remark 2.39.
In particular, if $P(\sigma)$ satisfies the polytropic equations of state, then $\sigma A^{\prime}(\sigma)=\gamma A(\sigma)=\frac{K\gamma}{\gamma-1}\sigma^{\gamma}$. Thanks to Lemma 2.37, we have $E_{0}(\sigma)=(4-3\gamma)U(\sigma)=(4-3\gamma)\int_{\mathbb{R}^{3}}A(\sigma)\,dx$, and $G(\sigma,\sigma)=(6\gamma-6)U(\sigma)$. Actually, inspired by Lieb and Yau [42, Remark after Lemma 6], we claim
$G(\sigma,\sigma)=(6\gamma-6)U(\sigma)$
for minimizers using the scaling relations similar to those in Section 3. That is, (with an abuse of notations) let $\sigma_{\lambda}(x)=\lambda^{3}\sigma(\lambda x)$ we have $\int_{\mathbb{R}^{3}}\sigma_{\lambda}\,dx=\int_{\mathbb{R}^{3}}\sigma\,dx$ thus $E_{0}(\sigma)\leq E_{0}\left(\sigma_{\lambda}\right)$, $\left.\frac{d}{d\lambda}E_{0}\left(\sigma_{\lambda}\right)\right|_{\lambda=1}=0$. Moreover, $E_{0}\left(\sigma_{\lambda}\right)=\lambda^{3\gamma-3}U(\sigma)-\frac{\lambda}{2}G(\sigma,\sigma)$, then we have $\left.\frac{d}{d\lambda}E_{0}\left(\sigma_{\lambda}\right)\right|_{\lambda=1}=(3\gamma-3)U(\sigma)-\frac{G(\rho,\rho)}{2}=0$, and hence the claim follows.
In order to obtain uniqueness result for minimizers, Lieb and Yau [42, Lemma 11 and Remark following] use a proof by contradiction. They said the only property of $A(s)$ used was the convexity of $g$ (also see [46, Section 3]), where $g(s)=4A(s)-3A^{\prime}(s)s$. However, a closer look reveals that the proof needs the strict inequality $g^{\prime\prime}(s)<0$ on some interval, whereas convexity only gives $g^{\prime\prime}(s)\leq 0$. Nevertheless, it turns out we indeed can show that the convexity of $g$, together with the specific structure of $A$, implies $g^{\prime\prime}<0$ holds true in a small interval, which is already enough to make a contradiction in Proposition 2.42.
In the following lemma, instead of considering $g$, we consider $f(s):=A^{\prime}\left(s^{3}\right)$ and one can check $g^{\prime\prime}(s)<0$ is equivalent to $f^{\prime\prime}(s)>0$ when $s>0$.
Lemma 2.40 (Positive Second Order Derivative of $A^{\prime}\left(s^{3}\right)$).
If $P(s)$ satisfies (F1)(F2)(F3)(F4), $A(s)\in C^{3}\left(\mathbb{R}^{+}\right)$(that is $A(s)$ has third order continuous derivative when $s>0)$, and $f(s):=A^{\prime}\left(s^{3}\right)$ is convex. Then $\forall\delta>0,\exists s_{\delta}\in(0,\delta)$, such that $f^{\prime\prime}\left(s_{\delta}\right)>0$.
Proof.
Since $P^{\prime}(s)=sA^{\prime\prime}(s)$, we know $A(s)\in C^{3}\left(\mathbb{R}^{+}\right)$ is equivalent to $P(s)\in C^{2}\left(\mathbb{R}^{+}\right)$. Notice $f^{\prime}(s)=3s^{2}A^{\prime\prime}\left(s^{3}\right)=\frac{3P^{\prime}\left(s^{3}\right)}{s}$, and $f^{\prime\prime}(s)=\frac{9s^{2}P^{\prime\prime}\left(s^{3}\right)\cdot s-3P^{\prime}\left(s^{3}\right)}{s^{2}}$. Since $f(s)$ is convex, $f^{\prime\prime}(s)=\frac{9s^{2}P^{\prime\prime}\left(s^{3}\right)\cdot s-3P^{\prime}\left(s^{3}\right)}{s^{2}}\geq 0$. If the conclusion does not hold, then $\exists\delta_{0}>0$, such that $\forall s\in\left(0,\delta_{0}\right),f^{\prime\prime}(s)=0$, which implies $3tP^{\prime\prime}(t)-P^{\prime}(t)=0$ for all $t\in\left(0,\delta_{0}^{3}\right)$, then $P^{\prime}(t)=Ct^{\frac{1}{3}}$ for some $C$ when $t\in\left(0,\delta_{0}^{3}\right)$. Then $\lim\limits_{t\rightarrow 0}\frac{P^{\prime}(t)}{t^{\frac{1}{3}}}=C$. Since we already know $P(t)$ satisfies (F2), that is $\lim\limits_{t\rightarrow 0}\frac{P(t)}{t^{\frac{4}{3}}}=0$, then $P\rightarrow 0$ as $t\rightarrow 0$. Then we can apply L’Hôpital’s rule [52, Theorem 5.13] and obtain
$0=\lim\limits_{t\rightarrow 0}\frac{P(t)}{t^{\frac{4}{3}}}=\lim\limits_{t\rightarrow 0}\frac{3}{4}\cdot\frac{P^{\prime}(t)}{t^{\frac{1}{3}}}=\frac{3C}{4}$
which means $C=0$. Then $P^{\prime}(t)=Ct^{\frac{1}{3}}=0$ when $t\in\left(0,\delta_{0}^{3}\right)$, notice again by ((F2)) $\lim\limits_{t\rightarrow 0}\frac{P(t)}{t^{\frac{4}{3}}}=0$, hence $P(t)=0$. It contradicts the assumption (F1) that $P(t)$ is strictly increasing for $t>0$. Therefore, we know $\forall\delta>0,\exists s_{\delta}\in(0,\delta)$, such that $f^{\prime\prime}\left(s_{\delta}\right)>0$. ∎
Remark 2.41.
Since $f^{\prime\prime}$ is continuous, by Lemma 2.40 we know $\forall\delta>0$, $\exists\left(s_{1},s_{2}\right)\subset(0,2\delta)$ such that $f^{\prime\prime}(s)>0$ if $s\in\left(s_{1},s_{2}\right)$.
Proposition 2.42 (Uniqueness of Minimizer in Non-rotating Cases [42] [46]).
If $P(s)$ satisfies (F1)(F2)(F3)(F4), $A(s)\in C^{3}\left(\mathbb{R}^{+}\right)$, and $A^{\prime}\left(s^{3}\right)$ is convex, then the minimizer of $E_{0}(\rho)$ is unique up to translation. In particular, if $P(s)$ satisfies the polytropic equations of state with index $\gamma>\frac{4}{3}$, the minimizer of $E_{0}(\rho)$ is unique up to translation.
Proof.
We provide a proof outline here. One can check the proofs in [42] and references therein, to understand some step(s) omitted in the following.
Let $f(s)=A^{\prime}\left(s^{3}\right)$, and $g(s)=4A(s)-3A^{\prime}(s)s$, by direct computation we obtain $f^{\prime\prime}(s)=6sA^{\prime\prime}\left(s^{3}\right)+9A^{\prime\prime\prime}\left(s^{3}\right)s^{4}$, and $g^{\prime\prime}(s)=-2A^{\prime\prime}(s)-3sA^{\prime\prime\prime}(s)$ when $s>0$. (Notice $A(s)\in C^{3}\left(\mathbb{R}^{+}\right)$ guarantees those derivatives exist and are continuous.) When $s>0$, $f$ is convex then $f^{\prime\prime}\geq 0$ implies $g^{\prime\prime}\leq 0$, then $g$ is concave. Therefore, we have $\frac{g(a)-g(b)}{a-b}\left\{\begin{array}[]{l}\leq g^{\prime}(b),b<a\\ \geq g^{\prime}(b),b>a\end{array}\right$., therefore, $\forall a,b\geq 0$, we have $g(a)-g(b)\leq g^{\prime}(b)(a-b)$.
If $m=0$, then minimizer $\sigma_{m}=0$ is unique. If $m>0$, without loss of generality, we assume $m=$ 1. Given two minimizers $\sigma$ and $\widetilde{\sigma}$, thanks to Theorem 2.6, after translation, we can assume $\sigma$ and $\widetilde{\sigma}$ are continuous, spherically symmetric, centered at 0, differentiable when they are positive, and radially decreasing. Then we know $\sigma(0)>0,\widetilde{\sigma}(0)>0$, and $\nabla\sigma(0)=\nabla\widetilde{\sigma}(0)=0$ (due to being symmetric and differentiable at 0). If $\sigma(0)=\widetilde{\sigma}(0)$, due to (30) (36) we have $\Theta_{\sigma}(0)=\Theta_{\widetilde{\sigma}}(0)$, and $\dot{\Theta}_{\sigma}(0)=\dot{\Theta}_{\widetilde{\sigma}}(0)$. By the uniqueness theorem of ODE, $\Theta_{\sigma}=\Theta_{\widetilde{\sigma}}$ where they are positive. Due to (36) we obtain $\sigma=\widetilde{\sigma}$. If $\sigma(0)\neq\widetilde{\sigma}(0)$, without loss of generality, we assume $\sigma(0)>\widetilde{\sigma}(0)>0$. Let $R>0$ and $\widetilde{R}>0$ be the radii of their supports. By Lemma 2.35, $m(r)>\widetilde{m}(r)$ for all $0<r<\max\{R,\widetilde{R}\}$. Then $R\leq\widetilde{R}$; otherwise $m=\widetilde{m}(\widetilde{R})<m(\widetilde{R})<m$, which is a contradiction. Then since $\sigma$ and $\widetilde{\sigma}$ are minimizers, thanks to Lemma 2.37, use spherical coordinate transformation then we have
[…]
$=\int_{\mathbb{R}^{3}}g(\sigma)-g(\widetilde{\sigma})\,dx$
$=4\pi\int_{0}^{\widetilde{R}}(g(\sigma(r))-g(\widetilde{\sigma}(r)))r^{2}dr$
$\leq 4\pi\int_{0}^{\widetilde{R}}g^{\prime}(\widetilde{\sigma}(r))(\sigma(r)-\widetilde{\sigma}(r))r^{2}dr$
[…]
Corollary 2.43.
Let $\sigma_{m}$ be the minimizer for mass $m$, then the mass $m$ can be viewed as a strictly increasing function of minimizers’ central density.
Proof.
If $\sigma_{m}$ and $\sigma_{m^{\prime}}$ are minimizers for mass $m$ and $m^{\prime}$ respectively, thanks to Theorem 2.6 (iii) we can assume they are spherically symmetric and radially decreasing, and $\sigma_{m}(0)$ and $\sigma_{m^{\prime}}(0)$ are their central densities. If $\sigma_{m}(0)>\sigma_{m^{\prime}}(0)$, due to Lemma 2.35 we know $m\geq m^{\prime}$. But Proposition 2.42 tells us that $m=m^{\prime}$ is impossible unless $\sigma_{m}=\sigma_{m^{\prime}}$. Thus, we have $m>m^{\prime}$. ∎
Now we know the minimizer satisfies (EL) and it is unique. A natural question is: given a radial solution $\sigma_{(\alpha)}$ to (EL) with central density (initial data) $\sigma_{(\alpha)}(0)=\alpha$, is it a minimizer? It turns out the answer is yes, but we first need to discuss more about Remark 2.34 to tie together minimizers, radial solutions to (EL) and positive radial solutions to (37). We first review a lemma in [42] (but for simplicity, here we assume the pressure satisfies the polytropic equations of state).
Lemma 2.44 (Compact Support of $\Theta$ [42, Lemma 8]).
Assume the pressure satisfies the polytropic equations of state with index $\gamma>\frac{3}{2}$. Let $\Theta$ be the radial solution(i.e., $\Theta$ is spherically symmetric) to (37) obtained by integrating outwards from $r=$ 0 with the initial conditions $\Theta(0)=\beta>0$ and $\dot{\Theta}(0)=0$. Then there is some $R(\beta)\in(0,\infty)$ such that $\Theta(R(\beta))=0$. For $r<R(\beta),\dot{\Theta}(r)<0$. In particular, $\Theta(r)$ is decreasing up to $R(\beta)$.
Proof.
$\Theta(0)=\beta>0$ implies $R(\beta)\neq 0$ if $R(\beta)$ exists. Set $\rho=\left(A^{\prime}\right)^{-1}(\Theta)$, by Shell Theorem, or the arguments similar as (26) and (28) we have $\Theta^{\prime}(r)=-4\pi\int_{0}^{r}\frac{t^{2}}{r^{2}}\rho(t)dt$, from which we get $\Theta^{\prime}(r)<0$ for $r<R(\beta)$. Integrate it again and use Fubini’s theorem, we have
$\Theta(r)=\beta-4\pi\int_{0}^{r}\left(\frac{1}{t}-\frac{1}{r}\right)t^{2}\rho(t)dt$
Suppose $R(\beta)=\infty$, set $g(t)=t\rho(t)$, we could obtain the result that $g\in L^{1}([0,\infty),dr)$ [42, Lemma 8]. However, one can also show $\Theta(r)\geq\frac{c}{r}$ for large $r$ [42, Lemma 8]. Then for $r$ large enough we have
$\rho=\left(A^{\prime}\right)^{-1}(\Theta)=C\Theta^{\frac{1}{\gamma-1}}\geq cr^{-\frac{1}{\gamma-1}}$
Notice $\gamma>\frac{3}{2}$ implies $1-\frac{1}{\gamma-1}>-1$, thus $\int_{0}^{\infty}g(r)dr=\int_{0}^{\infty}r\rho(r)dr=\infty$, which make a contradiction. For more details one can refer to [42, Lemma 8 and Proposition 9]. ∎
Remark 2.45.
Actually $\Theta^{\prime}(r)=-4\pi\int_{0}^{r}\frac{t^{2}}{r^{2}}\rho(t)dt$ implies $\Theta^{\prime}(r)<0$ when $\Theta$ is positive. Therefore, by contradiction arguments we can show $\Theta(r)\leq 0$ when $r\geq R(\beta)$, where $R(\beta)$ is given in Lemma 2.44.
Due to Remark 2.34, we know given a radial solution $\sigma_{(\alpha)}$ to (EL) with compact support, we can find a positive radial solution $\Theta_{\sigma_{(\alpha)}}$ to (37). On the other hand, Lemma 2.44 and Remark 2.45 tell us a radial solution $\Theta$ to (37) with positive initial data is only positive up to $r=R(\beta)$. Let
$\sigma_{\Theta}(r)=\left\{\begin{array}[]{r}\frac{-\Delta\Theta}{4\pi}\text{, }r\leq R(\beta)\\ 0\text{, }r>R(\beta)\end{array}\right$.
Then one can see $\sigma_{\Theta}$ satisfies (EL) with $\lambda=\Theta(0)-V_{\sigma_{\Theta}}(0)$.
We claim any radial solution $\sigma_{(\alpha)}$ to (EL) has compact support. In fact, due to (35) we know $\sigma_{(\alpha)}$ is decreasing when it is positive. If its support is unbounded, then $\sigma_{(\alpha)}(r)>0$ for all $r\geq 0$. Then we can construct a solution $\Theta_{\sigma_{(\alpha)}}$ to (37) with $\sigma_{(\alpha)}=\left(A^{\prime}\right)^{-1}\left(\Theta_{\sigma_{(\alpha)}}\right)$ (one can also see (36)), which implies $\Theta_{\sigma_{(\alpha)}}$ also has unbounded support. But it contracts Remark 2.45.
Therefore, we can actually remove the compact support assumption at the beginning in this remark that “a radial solution $\sigma_{(\alpha)}$ to (EL) with compact support”, and show the solutions to (EL) indeed correspond one-to-one with the positive solutions to (37).
Moreover, from $\lambda=\Theta(0)-V_{\sigma_{\Theta}}(0)$ above, one can see the choice of $\lambda$ in (EL) is not arbitrary but is related to the choice of $\Theta(0)$, which in turn is related to the central density $\sigma_{(\alpha)}(0)=\alpha$ (Recall $\sigma_{(\alpha)}=\left(A^{\prime}\right)^{-1}\left(\Theta_{\sigma_{(\alpha)}}\right)$).
Based on the above observation, we can now establish the equivalent relation between radial solutions to (EL) (with specific central densities) and minimizers:
Proposition 2.46 (Equivalence between (EL) Solutions and Minimizers [42, Lemma 12]).
Let $\sigma_{(\alpha)}$ be the unique nonnegative radial solution of (EL) with central density $\sigma(0)=\alpha$, then $\sigma_{(\alpha)}$ is the unique minimizer for $E_{0}(\rho)$ among $mR(\mathbb{R}^{3})$, where $m=\int_{\mathbb{R}^{3}}\sigma_{(\alpha)}\,dx$. That is, all radial solutions of (EL) parametrized by their central density are in fact minimizers of $E_{0}(\rho)$ among $mR(\mathbb{R}^{3})$ for some mass $m$.
Proof.
Here we provide a proof outline. One can refer to [42, Lemma 12] to understand some step(s) omitted here. We first note that the one-to-one correspondence between a central density of zero and a total mass of zero is trivial, as both correspond to the vacuum (zero function). Hence we can now turn to the case where mass is positive.
The idea of Lieb and Yau is: let $G=(0,\infty)$, and
[…]
For each $m\in G$, we know by Theorem 2.6 and Proposition 2.42 there is a unique minimizer $\rho_{m}$, and a unique central density $\alpha_{m}$ (see also Corollary 2.43). We let $\Gamma:G\rightarrow D$ denote this map from $m\in G$ to $\alpha_{m}\in D$.
Then in [42, Lemma 12], one can show $\Gamma$ (or $\Gamma^{-1}$) is a homeomorphism between $G$ and $D$, and $D$ turns out to be closed in $\mathbb{R}^{+}$. $G$ is connected, so is $D$, hence $D$ is a closed interval in $\mathbb{R}^{+}$. Since $\Gamma^{-1}$ is a homeomorphism, then $D$ can only be $\mathbb{R}^{+}$. ∎
Remark 2.47.
If assumption (F3) is replaced by (F3’), then $G$ in the proof of Proposition 2.46 should be $(0,m(K))$, where $m(K)$ is the Chandrasekhar mass mentioned in Remark 2.13.
3 Scaling Relations between Stars with Different Mass
The goal of this section is to provide a description of scaling method and scaling relations between solutions with different mass, which can help us to find more quantitative properties. We have been using those results in the previous arguments and will continue to use them in the following. We assume the polytropic equation of state $P(\rho)=K\rho^{\gamma}$ with index $\gamma>\frac{4}{3}$ holds true from this section.
3.1 Scaling relations between solutions to partial differential equations
It is known that the Navier-Stokes equations (and many other PDEs) have a natural scaling. The (incompressible) Navier-Stokes equations have the form:
$\left\{\begin{array}[]{l}\partial_{t}u+(u\cdot\nabla)u+\nabla\pi=v\Delta u\\ \operatorname{div}(u)=0\end{array}\right$. (NS)
where $\pi$ is the pressure and $v>0$ is the viscosity coefficient. For simplicity we assume $v=1$ is constant. Formally, we take the divergence of the first equation and use the second equation to eliminate some terms and obtain $\Delta\pi=-\operatorname{div}((u\cdot\nabla)u)$, i.e.
$\operatorname{div}((u\cdot\nabla)u)+\operatorname{div}(\nabla\pi)=0$
Then we can introduce Leray projection operator $\mathbb{P}=Id+\nabla(-\Delta)^{-1}\operatorname{div}$ [18] to get rid of the pressure term from the first equation and obtain:
$\left\{\begin{array}[]{l}\partial_{t}u+\mathbb{P}((u\cdot\nabla)u)=\Delta u\\ \operatorname{div}(u)=0\end{array}\right$. (NS’)
For any function $f(t,x)$, we denote $f_{\lambda}(t,x):=\lambda f\left(\lambda^{\alpha}t,\lambda^{\beta}x\right)$. Let a function $u(t,x)$ be a solution to (NS’), notice that $\mathbb{P}\left(f_{\lambda}\right)=(\mathbb{P}(f))_{\lambda}$, then by direct computation we have
$\partial_{t}u_{\lambda}+\mathbb{P}\left(\left(u_{\lambda}\cdot\nabla\right)u_{\lambda}\right)-\Delta u_{\lambda}=\lambda^{\alpha}\left(\partial_{t}u\right)_{\lambda}+\lambda^{1+\beta}(\mathbb{P}((u\cdot\nabla)u))_{\lambda}-\lambda^{2\beta}(\Delta u)_{\lambda}$
Therefore, when $\alpha=2$, $\beta=1$, $u_{\lambda}(t,x)=\lambda u\left(\lambda^{2}t,\lambda x\right)$ is also a solution to (NS’). Actually $\forall c\in\mathbb{R}$, $cu\left(\lambda^{2}t,\lambda x\right)$ is a solution since (NS’) is linear with respect to $c$.
However, when we consider the reduced Euler-Poisson equations (EP’) with $\omega=0$, or the corresponding Euler-Lagrange equations (EL) with $J=0$ (non-rotating case), since the pressure $P(\rho)=K\rho^{\gamma}$ or the function $A^{\prime}(\rho)=\frac{K}{\gamma-1}\rho^{\gamma}$ is not linear and there is an extra term $V_{\rho},c$ may not be chosen arbitrarily. With an abuse of notation, we assume $g_{A,B}(x)=Ag(Bx)$. Let a function $\sigma(x)$ be a solution to (EL), i.e., $A^{\prime}(\sigma(x))=\left[V_{\sigma}(x)+\lambda\right]_{+}$, then we have
$A^{\prime}\left(\sigma_{\widetilde{A},B}(x)\right)=\frac{K\gamma}{\gamma-1}\left(\sigma_{\widetilde{A},B}(x)\right)^{\gamma-1}=\widetilde{A}^{\gamma-1}A^{\prime}(\sigma(Bx))$
$V_{\sigma_{A,B}}(x)=\int_{\mathbb{R}^{3}}\frac{\widetilde{A}\sigma(By)}{|x-y|}\,dy=\widetilde{A}B^{-2}\int_{\mathbb{R}^{3}}\frac{\sigma(By)}{|Bx-By|}\,d(By)=\widetilde{A}B^{-2}V_{\sigma}(Bx)$
Hence if $\widetilde{A}^{\gamma-1}=\widetilde{A}B^{-2}$ (i.e., $B=\widetilde{A}^{\frac{2-\gamma}{2}}$), we will have $\sigma_{\widetilde{A},B}$ is also a solution to (EL):
$A^{\prime}\left(\sigma_{\widetilde{A},B}(x)\right)=\left[V_{\sigma_{\bar{A},B}}(x)+\lambda_{\widetilde{A},B}\right]_{+}$
with $\lambda_{\widetilde{A},B}=\widetilde{A}B^{-2}$.
Moreover, if we know $g$ has mass $m$ (again without loss of generality we assume $m\neq 0$ unless otherwise specified) and hope $g_{A,B}$ has mass 1. Then $\int_{\mathbb{R}^{3}}g_{A,B}\,dx=AB^{-3}$ $\int_{\mathbb{R}^{3}}g(Bx)\,d(Bx)=AB^{-3}m=1$. Therefore, we can solve $A=m^{-\frac{2}{3\gamma-4}},B=m^{\frac{\gamma-2}{3\gamma-4}}$.
Theorem 3.1 (Relations between Solutions to Equations with Different Mass).
Let a function $g(x)$ with mass $m$ be a solution to (EL) with $J=0$ and Lagrange multiplier $\lambda$, or to (EP’) with $\omega=0$, then $g_{A,B}(x)$ is a solution to (EL) with $J=0$ and Lagrange multiplier $\lambda_{A,B}$, or to (EP’) with $\omega=0$, where $A=m^{-\frac{2}{3\gamma-4}},B=m^{\frac{\gamma-2}{3\gamma-4}}$. Moreover, $g_{A,B}(x)$ has mass 1.
Proof.
For the case of Euler-Lagrange equations (EL), one can see the arguments above. For the case of the reduced Euler-Poisson equations (EP’), notice (EP’) becomes
$(\nabla P(g(x)))-g(x)\left(\nabla V_{g}(x)\right)=0$
We also notice
$\left(\nabla P\left(g_{A,B}(x)\right)\right)-g_{A,B}(x)\left(\nabla V_{g_{A,B}}(x)\right)=A^{\gamma}B\left(\nabla_{Bx}P(g(Bx))\right)-A^{2}B^{-1}g(Bx)\left(\nabla_{Bx}V_{g}(Bx)\right)$
thus the arguments are essentially the same since we have ${A}^{\gamma-2}B^{2}=1$. ∎
Therefore, to solve (EL), we can also assume a solution has mass 1. Notice the radial solution with certain mass is actually a minimizer and unique (see Proposition 2.42 and Proposition 2.46) and vice versa. It can help us to discuss more properties and relations between minimizers with different mass in the next subsection.
3.2 Scaling relations between minimizers of $E_{0}(\rho)$ with different mass
In this subsection we will show that given a minimizer, the corresponding scaling density can also be a minimizer, and then we will discuss more quantum properties.
Recall (5) that $E_{0}(\rho)=U(\rho)-\frac{G(\rho,\rho)}{2}=\int_{\mathbb{R}^{3}}A(\rho)\,dx-\frac{1}{2}\iint_{{\mathbb{R}^{3}}\times{\mathbb{R}^{3}}}\frac{\rho(x)\rho(y)}{|x-y|}\,dxdy$, where $A(\rho)=\frac{K}{\gamma-1}\rho^{\gamma}$.
Let $\sigma$ be in the admissible class (4), then we have
$U\left(\sigma_{\widetilde{A},B}\right)$ $=\int_{\mathbb{R}^{3}}A\left(\sigma_{\widetilde{A},B}\right)\,dx$
$=\int_{\mathbb{R}^{3}}\frac{K}{\gamma-1}(\widetilde{A}\sigma(Bx))^{\gamma}\,dx$
$=\frac{\widetilde{A}^{\gamma}}{B^{3}}\int_{\mathbb{R}^{3}}\frac{K}{\gamma-1}(\sigma(Bx))^{\gamma}\,d(Bx)$
$=\widetilde{A}^{\gamma}B^{-3}U(\sigma)$
$G\left(\sigma_{\widetilde{A},B},\sigma_{\widetilde{A},B}\right)$ $=\iint_{{\mathbb{R}^{3}}\times{\mathbb{R}^{3}}}\frac{\sigma_{\widetilde{A},B}(x)\sigma_{\widetilde{A},B}(y)}{|x-y|}\,dxdy$
$=\widetilde{A}^{2}B^{-5}\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\frac{\sigma(Bx)\sigma(By)}{|Bx-By|}\,d(Bx)\,d(By)$
$=\widetilde{A}^{2}B^{-5}G(\sigma,\sigma)$
$\int_{\mathbb{R}^{3}}\sigma_{\widetilde{A},B}\,dx$ $=\widetilde{A}B^{-3}\int_{\mathbb{R}^{3}}\sigma(Bx)\,d(Bx)$
$=\widetilde{A}B^{-3}\int_{\mathbb{R}^{3}}\sigma\,dx$
Assume $\sigma_{m}$ is a minimizer of $E_{0}(\rho)$ on $m{R}\left(\mathbb{R}^{3}\right)$, if $A^{\gamma}B^{-3}=A^{2}B^{-5}$ (i.e., $B=A^{\frac{2-\gamma}{2}}$), we will have $\left(\sigma_{m}\right)_{A,B}$ is a minimizer of $A^{\gamma}B^{-3}E_{0}(\rho)$ on $AB^{-3}m{R}\left(\mathbb{R}^{3}\right)$ (one can show it by contradiction arguments), and thereby a minimizer of $E_{0}(\rho)$ on $AB^{-3}m{R}\left(\mathbb{R}^{3}\right)$. If we hope $\left(\sigma_{m}\right)_{A,B}$ has mass $1$, then $AB^{-3}m=1$. Therefore, we can solve $A=m^{-\frac{2}{3\gamma-4}},B=m^{\frac{\gamma-2}{3\gamma-4}}$, which are compatible with the results in subsection 3.1 as one can expect.
Theorem 3.2 (Relations between Minimizers and Minimal Energies with Different Mass).
$E_{0}(\rho)$ allows a unique minimizer on $m{R}\left(\mathbb{R}^{3}\right)$ up to translation. Let $\sigma$ be the minimizer with mass 1 and the corresponding minimal energy is $e_{0}=E_{0}(\sigma)$, then any other minimizer with mass $m$ can be represented as $\sigma_{m}(x)=\frac{1}{A}\sigma\left(\frac{1}{B}x\right)$, where $A=m^{-\frac{2}{3\gamma-4}},B=m^{\frac{\gamma-2}{3\gamma-4}}$, and the corresponding minimal energy is $e_{0}(m)=E_{0}\left(\sigma_{m}\right)=m^{\frac{5\gamma-6}{3\gamma-4}}e_{0}$.
Proof.
The uniqueness result comes from Proposition 2.42. The relation between $\sigma_{m}$ and $\sigma$ can be seen in the arguments above. Notice $A^{-\gamma}B^{3}=m^{\frac{5\gamma-6}{3\gamma-4}}$, thus we have $e_{0}(m)=m^{\frac{5\gamma-6}{3\gamma-4}}e_{0}$. ∎
Remark 3.3.
Thanks to Theorem 2.6, we know $\|\sigma\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\leq C_{1}$ and spt $\sigma$ is contained in a ball of radius $R_{1}$, therefore, $\left\|\sigma_{m}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\leq C_{m}=\frac{C_{1}}{A}$ and spt $\sigma_{m}$ is contained in a ball of radius $R_{m}=BR_{1}$. Since $A=m^{-\frac{2}{3\gamma-4}},B=m^{\frac{\gamma-2}{3\gamma-4}}$, if we further assume $\gamma>2$, we know $\lim\limits_{m\rightarrow 0}A=+\infty$ and $\lim\limits_{m\rightarrow 0}B=0$, thus $\left\|\sigma_{m}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}$ and the size of $\sigma_{m}$’s support will go to 0 when $m\rightarrow 0$, with rates $m^{\frac{2}{3\gamma-4}}$ and $m^{\frac{\gamma-2}{3\gamma-4}}$ respectively. Notice $\left\|\sigma_{m}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}$ is actually the central density of $\sigma_{m}$, thus it gives the decay rate and strengthens Corollary 2.43. However, when $\gamma<2$, $\lim\limits_{m\rightarrow 0}B=\infty$, the size of $\sigma_{m}$’s support can go to $\infty$ when $m\rightarrow 0$ (flatten out or dispread). It coincides with a result mentioned in Lieb and Yau’s paper [42, Theorem 5], which says that the radius $R_{m}\rightarrow\infty$ as $m\rightarrow 0$. In their paper quantum mechanics (fermions case) is discussed.
Since we have the scaling relationship between energies, intuitively, one might guess whether a scaling relationship exists between the derivatives. Indeed, there is a relationship.
Proposition 3.4 (Relations of Variational Derivatives Between Densities with Different Mass).
Given $\rho_{m}\in mR(\mathbb{R}^{3})$ with $U(\rho_{m})<\infty$, then $E_{0}\left(\rho_{m}\right)$ is $P_{\infty}\left(\rho_{m}\right)$ differentiable at $\rho_{m}$ and the derivative at $\rho_{m}$ satisfies
$E_{0}^{\prime}\left(\rho_{m}\right)(x)=A^{-\gamma-1}E_{0}^{\prime}(\rho)\left(\frac{x}{B}\right)=m^{\frac{2\gamma-2}{3\gamma-4}}E_{0}^{\prime}(\rho)\left(\frac{1}{B}x\right)$
where $\rho(x)\coloneq\left(\rho_{m}\right)_{A,B}(x)=A\rho_{m}(Bx)\in R(\mathbb{R}^{3})$ with $U(\rho)<\infty$, $A=m^{-\frac{2}{3\gamma-4}}$, $B=m^{\frac{\gamma-2}{3\gamma-4}}$, $\rho$ has mass 1.
Proof.
The result $E_{0}\left(\rho_{m}\right)$ is $P_{\infty}\left(\rho_{m}\right)$-differentiable at $\rho_{m}$ comes from Lemma 2.9. Also easy to check $U(\rho)<\infty$ and has mass 1, and $E_{0}(\rho)$ is $P_{\infty}(\rho)$-differentiable at $\rho$. Given $\tau_{m}\in P_{\infty}\left(\rho_{m}\right)$, one can check $\tau:=\left(\tau_{m}\right)_{A,B}$ is in $P_{\infty}(\rho)$. By the definition of variational derivatives, one has
$\int_{\mathbb{R}^{3}}E_{0}^{\prime}\left(\rho_{m}\right)\tau_{m}\,dx$ $=\lim_{t\rightarrow\infty}\frac{E_{0}\left(\rho_{m}+t\tau_{m}\right)-E_{0}\left(\rho_{m}\right)}{t}$
$=\lim_{t\rightarrow\infty}\frac{E_{0}\left((\rho+t\tau)_{\frac{1}{A},\frac{1}{B}}\right)-E_{0}\left(\rho_{\frac{1}{A},\frac{1}{B}}\right)}{t}$
$=\lim_{t\rightarrow\infty}A^{-\gamma}B^{3}\frac{E_{0}(\rho+t\tau)-E_{0}(\rho)}{t}$
$=A^{-\gamma}B^{3}\int_{\mathbb{R}^{3}}E_{0}^{\prime}(\rho)\tau\,dx$
$=A^{-\gamma-1}\int_{\mathbb{R}^{3}}E_{0}^{\prime}(\rho)(x)\tau_{m}(Bx)\,d(Bx)$
$=A^{-\gamma-1}\int_{\mathbb{R}^{3}}E_{0}^{\prime}(\rho)\left(\frac{x}{B}\right)\tau_{m}(x)\,dx$
Since there are enough functions $\tau_{m}$ in $P_{\infty}\left(\rho_{m}\right)$, thus we have
$E_{0}^{\prime}\left(\rho_{m}\right)(x)=A^{-\gamma-1}E_{0}^{\prime}(\rho)\left(\frac{x}{B}\right)=m^{\frac{2\gamma-2}{3\gamma-4}}E_{0}^{\prime}(\rho)\left(\frac{1}{B}x\right)$
Remark 3.5.
Suppose $U(\rho_{m})<\infty$, there is another way to show the result. Let $\rho(x)=A\rho_{m}(Bx)$ with $A=m^{-\frac{2}{3\gamma-4}}$ and $B=m^{\frac{\gamma-2}{3\gamma-4}}$, we already know $E_{0}^{\prime}(\rho)=A^{\prime}(\rho(x))-V_{\rho}(x)$, and then one can also compute explicitly that
$A^{\prime}\left(\rho_{m}\right)(x)=A^{\prime}\left(\frac{1}{A}\rho\right)\left(\frac{1}{B}x\right)=\frac{1}{A^{\gamma-1}}A^{\prime}(\rho)\left(\frac{1}{B}x\right)$
$V_{\rho_{m}}(x)=\frac{B^{2}}{A}V_{\rho}\left(\frac{1}{B}x\right)=\frac{1}{A^{\gamma-1}}V_{\rho}\left(\frac{1}{B}x\right)$
and obtain again
$E_{0}^{\prime}\left(\rho_{m}\right)(x)=\frac{1}{A^{\gamma-1}}E_{0}^{\prime}(\rho)\left(\frac{1}{B}x\right)=m^{\frac{2\gamma-2}{3\gamma-4}}E_{0}^{\prime}(\rho)\left(\frac{1}{B}x\right)$
Appendix
Appendix A Properties of Sobolev Spaces
We recall some properties of Sobolev spaces as well as of $L^{p}$ spaces. Some of them will be used frequently in this paper.
Proposition A.1 (Hardy-Littlewood-Sobolev Inequality [5, Theorem 1.7]).
Let $1<p,r<\infty$ and $0<\alpha<n$ be such that $\frac{1}{p}+\frac{\alpha}{n}=\frac{1}{r}+1$. $\exists C_{p,\alpha,n}>0$, such that
$\left\||\cdot|^{-\alpha}*f\right\|_{L^{r}\left(\mathbb{R}^{n}\right)}\leq C_{p,\alpha,n}\|f\|_{L^{p}\left(\mathbb{R}^{n}\right)}$ (40)
By Hardy-Littlewood-Sobolev Inequality Proposition A.1, we can show the potential energy is bounded by density’s $L^{1}$ norm (density’s mass) and density’s $L^{\frac{4}{3}}$ norm.
Proposition A.2 (Bound of Potential Energy [4, Proposition 6]).
$\exists C>0$, if $\rho\in L^{1}\left(\mathbb{R}^{3}\right)\cap L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$, let $V_{\rho}(x)=$ $\int_{\mathbb{R}^{3}}\frac{\rho(y)}{|x-y|}\,dy$, then
$\left|\int_{\mathbb{R}^{3}}\rho\cdot V_{\rho}\,dx\right|\leq C\int_{\mathbb{R}^{3}}|\rho|^{\frac{4}{3}}\,dx\cdot\left(\int_{\mathbb{R}^{3}}|\rho|\,dx\right)^{\frac{2}{3}}$
Notice that Hardy-Littlewood-Sobolev Inequality Proposition A.1 can fail when $r=\infty$. To estimate the bound in $L^{\infty}$, we introduce the following proposition:
Proposition A.3 (Bound of Potential [4, Proposition 5] [17, Proposition B.1]).
Suppose $\rho\in L^{1}\left(\mathbb{R}^{3}\right)\cap L^{p}\left(\mathbb{R}^{3}\right)$. If $1<p\leq\frac{3}{2}$, then $\forall r\in\left(3,\frac{3p}{3-2p}\right)$, we have $V_{\rho}\in L^{r}\left(\mathbb{R}^{3}\right)$, and $\exists 0<b_{r}<1,0<c_{r}<1,C>0$, such that
$\left\|V_{\rho}\right\|_{L^{r}}\leq C\left(\|\rho\|_{L^{1}}^{b_{r}}\|\rho\|_{L^{p}}^{1-b_{r}}+\|\rho\|_{L^{1}}^{c_{r}}\|\rho\|_{L^{p}}^{1-c_{r}}\right)$ (41)
If $p>\frac{3}{2}$, then $V_{\rho}$ is bounded and continuous and satisfies (41) with $r=\infty$.
As discussed in [17, Appendix B], it turns out $h(x)=-\int_{\mathbb{R}^{3}}\frac{y_{j}\rho(x-y)}{|y|^{3}}\,dy$ is the weak derivative of $V_{\rho}(x)$. Moreover, for $p>3$, we have $\frac{\partial V_{\rho}}{\partial x_{j}}(x)=-\int_{\mathbb{R}^{3}}\frac{y_{j}\rho(x-y)}{|y|^{3}}\,dy$ not only in the distribution sense but also in the classical sense:
Proposition A.4 (Differentiability of Potential [4, Section 3] [17, Proposition B.3]).
If $\rho\in L^{1}\left(\mathbb{R}^{3}\right)\cap L^{p}\left(\mathbb{R}^{3}\right)$ for some $p>3$, then $V_{\rho}\in$ $W^{1,\infty}\left(\mathbb{R}^{3}\right)$ is continuously differentiable, and the weak derivative coincides with the classical one for all $x\in\mathbb{R}^{3}$.
To prove the low semicontinuity of the energy, we hope to connect weak convergence of densities and strong convergence of potentials.
Proposition A.5 (Compactness of Convolution (Potential) Operator with $3<q<\infty$).
Let $G(x)=\frac{1}{|x|}$, then $G\in L_{w}^{3}\left(\mathbb{R}^{N}\right)$. Let F=GBF=G*B, where $B$ is a bounded set in $L^{q}\left(\mathbb{R}^{N}\right)$ with $3<q<\infty$. Let $\Omega\subset\mathbb{R}^{3}$ be open, bounded and of class $C^{1}$, then $\left.F\right|_{\Omega}$ has compact closure in $C(\bar{\Omega})$. In particular, let $V_{\rho}(x)=\int_{\mathbb{R}^{3}}\frac{\rho(y)}{|x-y|}\,dy$, then for all $r\in[1,\infty]$, the map $\mathcal{V}:\rho\rightarrow V_{\rho}$ from $L^{q}(\Omega)$ to $L^{r}(\Omega)$ is compact.
Proof.
By Proposition A.4 we know $F\subset W^{1,\infty}\left(\mathbb{R}^{3}\right)$. Following its proof in [17] in detail one can see the norm $\|\cdot\|_{W^{1,\infty}}$ is bounded uniformly. The result comes from Rellich-Kondrachov Theorem [8, Theorem 9.16]. ∎
Actually, one can have a more general version:
Proposition A.6 (Compactness of Convolution (Potential) Operator with $q\geq 1$).
Let $G(x)=\frac{1}{|x|}$, then $G\in L_{w}^{3}\left(\mathbb{R}^{3}\right)$. Let F=GBF=G*B, where $B$ is bounded both in $L^{1}\left(\mathbb{R}^{3}\right)$ and $L^{q}\left(\mathbb{R}^{3}\right)$ with $q>1$. Let $\Omega\subset\mathbb{R}^{3}$ be open, bounded and of class $C^{1}$, then
  1. (1)
    if $1<q<\frac{3}{2}$, then $\forall 1\leq r<p$, where $\frac{1}{p}=\frac{1}{q}-\frac{2}{3},\left.F\right|_{\Omega}$ has compact closure in $L^{r}(\Omega)$;
  2. (2)
    if $q=\frac{3}{2}$, then $\forall 1\leq r<\infty,\left.F\right|_{\Omega}$ has compact closure in $L^{r}(\Omega)$;
  3. (3)
    if $q>\frac{3}{2}$, then $\left.F\right|_{\Omega}$ has compact closure in $C(\bar{\Omega})$.
Proof.
We first establish the boundedness of $F$ in $W^{1,s}$ for some appropriate $s$. Similar to the proof of [17, Proposition B.1] (cf. Proposition A.3 above), and noting that $K(y)=\frac{y_{j}}{|y|^{3}}$ is in $L^{p}_{\text{loc}}(\mathbb{R}^{3})$ for $p<3/2$ and belongs to $L^{p}(\mathbb{R}^{3}\setminus B_{R}(0))$ for $p>3/2$, we can show there exists an appropriate $s>0$, such that $\forall\rho\in B$, $V_{\rho}(x)$ and $\frac{\partial V_{\rho}}{\partial x_{j}}(x)=-\int_{\mathbb{R}^{3}}\frac{y_{j}\rho(x-y)}{|y|^{3}}\,dy$ are bounded in $L^{s}$. Since $B$ is bounded in $L^{1}$ and $L^{q}$, the bound of $V_{\rho}$ in $W^{1,s}$ is uniform for all $\rho\in B$. Once the boundedness of $F$ in $W^{1,s}$ is established, we apply the Rellich–Kondrachov theorem [8, Theorem 9.16] to conclude the proof.
Acknowledgments
The author is partially supported by the National Science Foundation grant DMS-2308208. This work was primarily carried out during the author’s Master’s studies at the University of Bonn. The author thanks Juan Velázquez and Dimitri Cobb for their continued advice and support since the author’s time in Bonn, as well as Christof Sparber and Mimi Dai for their comments and support during the author’s Ph.D. studies at the University of Illinois Chicago. Thanks also to Lorenzo Pompili, Shao Liu, Xiaopeng Cheng, Bernhard Kepka, Daniel Sánchez Simón del Pino for discussions, and to Théophile Dolmaire and other instructors. The author is grateful to his parents.